### v1

# On the dimension distortion under fractionally smooth mappings

Abstract.
We determine the extent to which continuous mappings in various Sobolev classes distort various dimensions, including the Hausdorff, the upper Minkowski (box-counting), and the upper intermediate dimensions. The intermediate and Minkowski dimension distortion results we obtain are novel already for various classes of fractionally smooth mappings between Euclidean spaces, extending the results of Hencl-Honzík [50, 49] and Huynh [51] to these dimensions. In addition, our work also generalizes the aforementioned results, as well as results of Kaufman [54] and Fraser-Tyson [31], to certain weighted Euclidean spaces and, more generally, to doubling metric measure spaces. As an application of our main result, we quantify a dimension distortion property of quasisymmetric mappings proved by Bishop-Hakobyan-Williams [16] for the intermediate dimension of non-Ahlfors regular subsets of the space.
2020 Mathematics Subject Classification:
Primary 30L10; Secondary 31E05, 28A80
1. Introduction
The class of Sobolev mappings has been famously an essential tool in the area of partial differential equations (PDEs) [25]. On the other hand, fractionally smooth Sobolev spaces provide a natural framework for problems where smoothness is intermediate between integer orders, acting as interpolation spaces in the context of functional analysis [1]. They arise naturally in the study of the fractional Laplacian, with applications on anomalous diffusion and jump processes, minimal surfaces, elliptic problems with measure data, and many other areas. We refer to [24] for an even more extensive list of applications and relevant references. In the past two and a half decades, there has been an increasing interest and need to extend this theory of (fractional) Sobolev and, more generally, Triebel–Lizorkin and Besov mappings to metric spaces. Applications of this endeavor include the development of the theory of PDEs [55], calculus of variations [8] and optimal transportation [9] on the non-smooth setting of fractal spaces. The theory of (fractional) Sobolev, Triebel–Lizorkin, and Besov mappings defined between metric spaces has been developed by many authors (see, for instance, [17, 38, 40, 41, 56, 62, 64, 57, 42, 63, 34] just to name a few), who have used different approaches to adjust the theory to different settings.
A question of broad interest has been to determine in what ways certain classes of mappings distort dimension notions. One of the earliest results in this direction is by Gehring-Väisälä [33], who gave quantitative bounds on how quasiconformal mappings, a special class of super-critical Sobolev mappings, change the Hausdorff dimension of a subset of $\mathbb{R}^{n}$. Kaufman later proved bounds for the distortion of the Hausdorff and Minkowski dimensions under general super-critical Sobolev mappings [54]. The study of dimension distortion has since been extended to sub-critical Sobolev mappings [48], fractionally smooth Sobolev mappings [50, 49, 51], to other dimension notions [22, 60, 19, 31], and to other settings, such as distortion by Sobolev and quasisymmetric mappings defined between metric spaces [13], [11], [12], [16]. However, in the non-Euclidean setting all results for Sobolev mappings are for the Hausdorff dimension, or cases where all dimensions coincide, with the distortion of the Minkowski dimension only recently settled by the second author [21]. Furthermore, there are currently no results on the dimension distortion under fractional Sobolev, Triebel–Lizorkin, or Besov mappings in the metric setting. In this manuscript, we address both of these open directions by considering the dimension distortion properties of compactly-Hölder mappings, a class that contains and unifies the notions of (fractional) Sobolev, Triebel–Lizorkin, and Besov mappings between metric spaces (see Section 2).
A modern approach to notions of fractal dimensions has been the introduction of dimension functions, which provide more information on the finer structure of spaces than typical dimension values. For instance, the Assouad spectrum, introduced by Fraser-Yu [32], is such a dimension function that naturally interpolates between the upper Minkowski and Assouad dimensions, based on the geometric properties of the space. We refer to [29] for an exposition on the topic, with a plethora of applications in areas such as number theory, probability, and functional analysis. A similar dimension function is the collection of (upper) intermediate dimensions, introduced by Falconer-Fraser-Kempton [28]. This notion interpolates between the Hausdorff and upper Minkowski dimensions, capturing finer geometric traits of the space not typically distinguished by the two extreme dimension values. Recent applications include towards the dimension theory of Brownian images [27], bi-Lipschitz classification of spaces [15], and the theory of orthogonal projections [30]. For a uniformly perfect metric space $X$ (see Section 2), the properties of $\theta$-intermediate dimensions of non-empty subsets $E$ of $X$, denoted by $\dim_{\theta}E$, were first established and studied by Banaji in [14]. In the same metric context, we establish intermediate dimension distortion bounds for compactly Hölder mappings, in the spirit of Gehring-Väisälä [33] and Kaufman [54]:
Theorem 1.1.
Suppose $(X,d_{X})$ is a doubling, uniformly perfect metric space and $(Y,d_{Y})$ is a uniformly perfect metric space. For $p\in(1,\infty)$ and $\alpha\in(0,\infty)$, if $f:X\rightarrow Y$ is $(p,\alpha)$-compactly Hölder and $E\subset X$ is bounded with $\dim_{\theta}E=d_{E}(\theta)$, then
(1.1) $\dim_{\theta}f(E)\leq\max\left\{\frac{pd_{E}(\theta)}{\alpha p+d_{E}(\theta)},d_{E}(\theta)\right\}.$
An immediate corollary is a similar dimension bound under Newtonian and quasisymmetric mappings. In particular, it follows by [21, Theorem 1.2] that a continuous mapping in the Newtonian-Sobolev class is also compactly Hölder for appropriate constants $p$ and $\alpha$, which also yields a similar inclusion for quasisymmetric mappings, under standard assumptions on $X$ and $Y$ (see [21, Corollary 1.3]).
Corollary 1.2.
  1. (i)
    Suppose $(X,d,\mu)$ is a proper, locally $Q$-homogeneous metric measure space supporting a $Q$-Poincaré inequality for some $Q\in(1,\infty)$, and $(Y,d_{Y})$ is an arbitrary uniformly perfect metric space. Let $f:X\rightarrow Y$ be a continuous mapping with an upper gradient $g\in L^{p}_{\scriptstyle{loc}}(X)$ with $p\in(Q,\infty)$. If $E\subset X$ is bounded with $\dim_{\theta}E=d_{E}(\theta)<Q$, then
    (1.2) $\dim_{\theta}f(E)\leq\frac{pd_{E}(\theta)}{p-Q+d_{E}(\theta)}<Q.$
  2. (ii)
    Suppose $Q\in(1,\infty)$ and $(X,d,\mu)$ is a proper and $Q$-Ahlfors regular metric measure space supporting a $p_{0}$-PI for $p_{0}\in(1,Q)$, and $(Y,d_{Y})$ is a $Q$-Ahlfors regular metric space. Let $f:X\rightarrow Y$ be an $\eta$-quasisymmetric homeomorphism. If $E\subset X$ is bounded with $\dim_{\theta}E=d_{E}(\theta)\in(0,Q)$, then
    (1.3) $0<\frac{(p-Q)d_{E}(\theta)}{p-d_{E}(\theta)}\leq\dim_{\theta}f(E)\leq\frac{pd_{E}(\theta)}{p-Q+d_{E}(\theta)}<Q,$
    where $p>Q$ only depends on $\eta(1),\eta^{-1}(1)$.
The Newtonian-Sobolev class constitutes one of the broader classes of Sobolev-type mappings between metric spaces (see Theorem 10.5.1 in [47]). Hence, the above corollary is a broad generalization of the result of Fraser-Tyson [31] and settles the intermediate dimension distortion problem on metric spaces, by providing a quantitative bound. A non-exhaustive list of spaces where the above result could be applied includes Carnot groups, Laakso spaces, Gromov hyperbolic groups and boundaries (see [47] Chapter 14 and references therein).
Bishop-Hakobyan-Williams [16] studied the quasisymmetric dimension distortion problem problem in the case where the input set $E$ is Ahlfors regular, which implies that all dimensions of $E$ coincide. Their motivation was the absolute continuity on lines property (ACL) that quasisymmetric mappings satisfy in the Euclidean setting. Their result provides a fundamental generalization of this fact in the metric measure spaces setting. In general, however, the Hausdorff, intermediate, and Minkowski dimensions of $E$ could all differ. In such a case, the results from [16] cannot be applied, while (1.3) provides quantitative bounds on $\dim_{\theta}f(E)$, which also recover [21, Theorem 1.2].
In order to reach the desired dimension distortion statement for (fractional) Sobolev, Triebel–Lizorkin, and Besov mappings, we need to ensure appropriate inclusions in the compactly Hölder class. This is achieved through embedding results of the metric Sobolev mappings of fractional smoothness in question, which generalize the first author’s work on real-valued functions [3, 7, 6] (see Section 3.2).
Theorem 1.3.
Suppose $(X,d_{X},\mu)$ is a locally $Q$-homogeneous metric measure space and $(Y,d_{Y})$ is an arbitrary metric space. Let $s\in(0,\infty)$, $p\in(Q/s,\infty)$, $q\in(0,\infty]$ and $f:X\rightarrow Y$ be a continuous mapping. If $f$ has a finite Hajłasz–Triebel–Lizorkin semi-norm $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}$, or a finite Hajłasz–Besov semi-norm $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}$ with $q\leq p$, then $f$ is (t,sQ/t)(t,s-Q/t)-compactly Hölder for all $t\in(Q/s,p)$. If $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}<\infty$ for $p<q<\infty$, then $f$ is (q,sQ/t)(q,s-Q/t)-compactly Hölder for all $t\in(Q/s,p)$. If, in addition to the assumptions above, $X$ is proper, and $X$ and $Y$ are uniformly perfect, then the following statements hold for any bounded set $E\subset X$ with $\dim_{\theta}E=d_{E}(\theta)<Q$.
  1. (i)
    If $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}<\infty$, or if $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}<\infty$ with $q\leq p$, then
    (1.4) $\dim_{\theta}f(E)\leq\max\left\{\displaystyle\frac{p\,d_{E}(\theta)}{sp-Q+d_{E}(\theta)},\;d_{E}(\theta)\right\}.$
  2. (ii)
    If $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}<\infty$ with $p<q<\infty$, then
    (1.5) $\dim_{\theta}f(E)\leq\max\left\{\frac{q\,d_{E}(\theta)}{(s-Q/p)q+d_{E}(\theta)},d_{E}(\theta)\right\}.$
The above theorem is in fact novel already in the Euclidean case $X=Y=\mathbb{R}^{n}$ with the usual measure and metric, as it is the first intermediate dimension distortion result for classical continuous Triebel–Lizorkin and Besov mappings. In addition, Theorem 1.3 also immediately yields dimension distortion bounds for continuous mappings whose fractional Hajłasz–Sobolev semi-norm $\|\cdot\|_{\dot{M}^{s,p}(X:Y)}$ is finite, since $\dot{M}^{s}_{p,\infty}(X:Y)=\dot{M}^{s,p}(X:Y)$; see Lemma 4.5. Furthermore, if $X,Y$ are not uniformly perfect, similar and simplified arguments that lead to Theorem 1.3 can be employed for $\dim_{0}E=\dim_{H}E$. Consequently, the work of Kaufman on Euclidean super-critical Sobolev mappings [54], the work of Hencl-Honzík on Euclidean Triebel-Lizorkin mappings [50], and the work of Huynh on Euclidean Besov mappings [51] are also recovered. In particular, we have the following unifying result for both the Hausdorff and the Minkowski dimensions.
Theorem 1.4.
Suppose $(X,d_{X},\mu)$ is a proper, locally $Q$-homogeneous metric measure space, $(Y,d_{Y})$ is an arbitrary metric space, and let $s$, $p$, $q$, and $f$ be as in Theorem 1.3.
  1. (i)
    If $E\subset X$ with $\dim_{H}E=a$, then
    (1.6) $\dim_{H}f(E)\leq\begin{cases}\displaystyle\max\left\{\frac{q\,a}{(s-Q/p)q+a},a\right\},&p<q<\infty\,\,\&\,\,\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}<\infty,\\[12.91663pt] \max\left\{\displaystyle\frac{p\,a}{sp-Q+a},\;a\right\},&p\geq q\,\,\&\,\,\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}<\infty,\\[12.91663pt] \max\left\{\displaystyle\frac{p\,a}{sp-Q+a},\;a\right\},&\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}<\infty.\end{cases}$
  2. (ii)
    If $E\subset X$ is bounded with $\dim_{B}E=a$, then (1.6) holds with $\dim_{H}f(E)$ replaced by $\dim_{B}f(E)$.
Although several approaches to defining Sobolev, Triebel–Lizorkin, and Besov spaces on metric spaces have been proposed over the years, a key advantage of the spaces considered in this work is that they support a robust theory without requiring the underlying metric space to be connected. Under additional assumptions on the metric measure space (e.g. Ahlfors regularity, Poincaré inequality), many of the proposed definitions are known to coincide and, hence, Theorems 1.3 and 1.4 can be used to provide dimension distortion bounds for an even broader class of Sobolev, Triebel–Lizorkin, and Besov spaces in that setting. We refer the interested reader to [5, 57, 34] and the references therein for more details.
Note that the bounds involving the Hausdorff and Minkowski dimensions in the above corollary are new for weighted Euclidean spaces, extending the respective results from [50, 49, 51] in that popular setting. For instance, if $\lambda_{n}$ is the $n$-Lebesgue measure, the conditions of Theorem 1.3 (and thus of Theorem 1.4) are satisfied by the weighted Euclidean metric measure space $(\mathbb{R}^{n},d_{\text{euc}},w\lambda_{n})$ for a wide variety of weights $w:\mathbb{R}^{n}\rightarrow[0,\infty]$, such as the class of Muckenhoupt weights (see Chapter 1 in [45]). These weights were introduced by Muckenhoupt [58] in order to characterize the boundedness of the Hardy-Littlewood maximal operator on weighted $L^{p}$ spaces, and have since established an active area within functional and harmonic analysis [36, 37]. More recently, certain Muckenhoupt weights have been associated with dimension theoretic characteristic of spaces through the “weak porosity” notion (see [10, 59]).
This paper is organized as follows. Section 2 includes the necessary background, and introduces an equivalent definition of the intermediate dimension using Hytönen-Kairem dyadic cubes. In Section 3 we prove the intermediate dimension distortion result under compactly Hölder maps (Theorem 1.1). Section 4 contains the proofs of Morrey-type embedding theorems, which we employ to show the Sobolev classes of fractional smoothness in question are contained in the appropriate compactly Hölder class and prove Theorem 1.3. Section 5 contains the proof for the Hausdorff and Minkowski dimension distortion under Hajłasz Triebel–Lizorkin Sobolev and Hajłasz Besov mappings (Theorem 1.4), as well as the proof of Corollary 1.2 for the distortion under Newtonian Sobolev and quasisymmetric mappings.
Acknowledgments.
Part of the project was completed during the first author’s visit to Knoxville and, thus, he wishes to thank the University of Tennessee for their hospitality.
2. Background
2.1. Metric spaces and dimensions.
Given two non-negative quantities $A$ and $B$, we write $A\lesssim B$ if there is a comparability constant $C=C(\lesssim)>0$ such that $A\leq CB$. Similarly, we write $A\gtrsim B$ if there is $C=C(\gtrsim)$ such that $A\geq B/C$. If $A\lesssim B$ and $A\gtrsim B$ we write $A\simeq B$.
Let $(X,d_{X})$ be a metric space. We often omit the subscript and write $d(x,y)$ for $x,y\in X$ if the space is understood. We denote the open ball centered at $x$ of radius $r>0$ by $B(x,r):=\{z\in X:\,\,d(x,z)<r\}.$ Given a ball $B=B(x,r)\subset X$, we denote by $\lambda B$ the ball $B(x,\lambda r)$, for $\lambda>0$. Given a non-empty set $U\subset X$, we denote by $|U|$ the diameter of $U$ in the metric of $X$. We also make the convention that all bounded sets we consider henceforth are non-empty, even if not explicitly stated, as all results trivially follow otherwise.
We say that $(X,d)$ is a doubling metric space if there is a doubling constant $C_{d}\geq 1$ such that for all $x\in X,r>0$, the smallest number of balls of radius $r$ needed to cover $B(x,2r)$ is at most $C_{d}$. Note that the doubling property implies that $X$ is separable.
We say that $(X,d)$ is a uniformly perfect metric space if there is $c_{u}\in(0,1)$ such that for every $x\in X$ and every $r<|X|$ there is a point $x^{\prime}\in B(x,r)\setminus B(x,c_{u}r)$. We say that $(X,d)$ is a $c_{u}$-uniformly perfect metric space if we need to emphasize the constant.
Let $E$ be a bounded subset of $X$. For $r>0$, denote by $N(E,r)$ the smallest number of sets of diameter at most $r$ needed to cover $E$. The (upper) Minkowski dimension of $E$ is defined as
${\overline{\dim_{B}}}(E)=\limsup_{r\to 0}\frac{\log N(E,r)}{\log(1/r)}.$
This notion is also known as upper box-counting dimension, which justifies the notation with the subscript ‘B’ typically used in the literature (see [26], [29]). We drop the adjective ‘upper’ and the bar notation throughout this paper as we will make no reference to the lower Minkowski dimension. If $X$ is a typical Euclidean space $\mathbb{R}^{n}$, an equivalent formulation for the Minkowski dimension is the following (see [28, Definition 1.1])
$\dim_{B}E=\inf\left\{\,d>0:\;\vbox{\hbox{$\begin{aligned} \\[-2.58334pt] &\forall\,\varepsilon>0\;\exists\,\delta_{\varepsilon}\in(0,1)\ \text{such that}\;\forall\,\delta\in(0,\delta_{\varepsilon})\;\text{there is}\\[1.29167pt] &\{U_{i}\}_{i}\ \text{cover of }E\ \text{with }\;|U_{i}|=\delta,\;\forall i,\;\text{and}\;\sum_{i}|U_{i}|^{d}<\varepsilon\end{aligned}$}}\right\}.$
On the other hand, if $|U_{i}|=\delta$ is replaced by $|U_{i}|\leq\delta$, the above definition would yield the Hausdorff dimension $\dim_{H}E$ of the set $E$. These representations motivated Falconer, Fraser, and Kempton to define the notion of intermediate dimensions [28], which is a dimension function, rather than a dimension value, geometrically interpolating between $\dim_{H}E$ and $\dim_{B}E$. While it was initially defined in [28] for the Euclidean setting, we state the definition for subsets $E$ of a uniformly perfect metric space $X$. For $\theta,\delta\in(0,1)$ we say that a cover $\{U_{i}\}_{i\in I}$ of $E$ is $\delta^{1/\theta}$-admissible if $\delta^{1/\theta}\leq|U_{i}|\leq\delta$, for all $i\in I$. The ($\theta$-upper-)intermediate dimension of $E$ is defined to be
$\dim_{\theta}E=\inf\left\{\,d>0:\;\vbox{\hbox{$\begin{aligned} \\[-2.58334pt] &\forall\,\varepsilon>0\;\exists\,\delta_{\varepsilon}\in(0,1)\ \text{such that}\;\forall\,\delta\in(0,\delta_{\varepsilon})\;\text{there is}\,\\[1.29167pt] &\{U_{i}\}_{i\in I}\;\delta^{1/\theta}\text{-admissible cover of }E\ \text{with }\;\;\sum_{i\in I}|U_{i}|^{d}<\varepsilon\end{aligned}$}}\right\}.$
Similarly to the Minkowski dimension convention, we make no mention to the lower intermediate dimension for the rest of the paper and, hence, we drop the adjective ‘upper’ in this case too. It should also be noted that in the uniformly perfect metric setting, if $c_{1},c_{2}>0$ are fixed constants and we slightly modify the notion $\delta$-admissible cover to require $c_{1}\delta\leq|U_{i}|\leq c_{1}\delta$ instead of equality, we similarly have
$\dim_{H}E=\dim_{0}E\qquad\text{and}\qquad\dim_{B}E=\dim_{1}E,$
with $\dim_{\theta}E$ being a continuous function of $\theta$ in $(0,1]$. We refer to [14] for a very interesting treatment of this, and other similar notions (generalized intermediate dimensions) in the metric setting.
On Euclidean spaces $X=\mathbb{R}^{n}$ with the usual metric one can use dyadic cubes instead of arbitrary sets of diameter between $\delta^{1/\theta}$ and $\delta$ to define the intermediate dimension (see [31]), and similarly for other dimension notions (see [26, 29]). On arbitrary metric spaces, however, there are various generalizations of dyadic cube constructions. One of the first manuscripts addressing this idea was by David [23], while one of the first explicit constructions of a system of dyadic cubes is due to Christ [18]. See also [2], [52], [53], [61], which is not an exhaustive list. The most fitting notion for our context is that of Hytönen and Kairema, which is enough to characterize various notions of dimension, including the Hausdorff, Assouad [20] and the Minkowski [21] dimensions.
Theorem A (Hytönen, Kairema [52]).
Suppose $(X,d)$ is a doubling metric space. Let $0<c_{0}\leq C_{0}<\infty$ and $b\in(0,1)$ with $12C_{0}b\leq c_{0}$. For any non-negative $k\in\mathbb{Z}$ and collection of points $\{z_{i}^{k}\}_{i\in I_{k}}$ with
(2.1) $d(z_{i}^{k},z_{j}^{k})\geq c_{0}b^{k},\,\,\,\,\text{for}\,\,i\neq j$
and
(2.2) $\min_{i}d(z_{i}^{k},x)<C_{0}b^{k},\,\,\,\,\text{for all}\,\,x\in X$
we can construct a collection of sets $\{Q_{i}^{k}\}_{i\in I_{k}}$ such that
  • (i)
    if $l\geq k$ then for any $i\in I_{k}$, $j\in I_{l}$ either $Q_{j}^{l}\subset Q_{i}^{k}$ or $Q_{j}^{l}\cap Q_{i}^{k}=\emptyset$,
  • (ii)
    $X$ is equal to the disjoint union $\bigcup\limits_{i\in I_{k}}Q_{i}^{k}$, for every $k\in\mathbb{N}$
  • (iii)
    $B(z_{i}^{k},c_{0}b^{k}/3)\subset Q_{i}^{k}\subset B(z_{i}^{k},2C_{0}b^{k})=:B(Q_{i}^{k})$ for every $k\in\mathbb{N}$,
  • (iv)
    if $l\geq k$ and $Q_{j}^{l}\subset Q_{i}^{k}$, then $B(Q_{j}^{l})\subset B(Q_{i}^{k})$.
For non-negative $k\in\mathbb{Z}$, we call the sets $Q_{i}^{k}$ from the construction of Theorem A ($b$-)dyadic cubes of level $k$ of $X$.
Fix $b$, $c_{0}$ and $C_{0}$ as in Theorem A. Moreover, for every non-negative $k\in\mathbb{Z}$ we fix a collection of points $\{z_{i}^{k}\}_{i\in I_{k}}$ and the corresponding collection of $b$-dyadic cubes $Q^{k}_{i}$. To see why such a collection of points exists, consider the covering $\{B(z,c_{0}b^{k}):z\in X\}$ of $X$ and apply the $5B$-covering lemma. By separability of $X$ and by choosing $c_{0}$ and $C_{0}$ so that $5c_{0}b^{k}<C_{0}b^{k}$, the existence of centers $\{z_{i}^{k}\}_{i\in I_{k}}$ is ensured. Given a doubling metric space $X$, we fix such a system of dyadic cubes for the rest of the paper. We also denote by $\mathcal{D}_{k}$ the collection of all cubes in the fixed system which are of level $k$.
Remark 2.1.
Note that given any $k$-level cube $Q$, there are at most a uniform number of cubes of level $k+1$ contained in $Q$, say $N_{d}\in\mathbb{N}$. This number $N_{d}$ solely depends on the doubling constant $C_{d}$ of $X$ and the constants $b$, $c_{0}$, $C_{0}$ of the dyadic system. Indeed, by Theorem A (iii), for every cube of level $k+1$ inside $Q$, there is a ball of radius $3^{-1}c_{0}b^{k+1}$ inside $B(Q)$, which is of radius $2C_{0}b^{k}$. Hence, by an application of the doubling property of $X$ we have at most
$N_{d}=C_{d}^{\left(\frac{2C_{0}b^{k}}{3^{-1}c_{0}b^{k+1}}\right)\log_{2}C_{d}}=C_{d}^{\left(\frac{2C_{0}}{3^{-1}c_{0}b}\right)\log_{2}C_{d}}$
such balls in $B(Q)$, which is the same bound on the number of $k+1$-level cubes inside $Q$.
We show that the intermediate dimension can be expressed using the dyadic cube systems from Theorem A. Note that a similar result in $\mathbb{R}^{n}$ was recently proved in [31] for the usual Euclidean dyadic cube system, although both the statement and the proof differ from those in the metric setting that are detailed below.
Proposition 2.2.
Let $E\subset X$ be a non-empty subset of a $c_{u}$-uniformly perfect $C_{d}$-doubling metric space $X$. For $\theta\in(0,1)$, the $\theta$-intermediate dimension of $E$ is the infimum of the set $A_{\theta}$ consisting of exponents $s>0$ for which for all $\varepsilon>0$ there exists $\delta_{\varepsilon}>0$ such that for all $\delta\in(0,\delta_{\varepsilon})$ there is a cover $\{Q_{i}\}_{i\in I_{\delta}}$ of $E$ by dyadic cubes $Q_{i}$ of level $k_{i}$ with
(2.3) $\frac{3}{c_{u}c_{0}}\,\delta^{1/\theta}\leq b^{k_{i}}\leq\frac{1}{4C_{0}}\delta,$
for all $i\in I_{\delta}$, and
$\sum_{i}|Q_{i}|^{s}<\varepsilon.$
Proof.
Note that if $Q_{i}$ is a cube of level $k_{i}$, then by Theorem A (iii) and uniform perfectness of $X$, we have
$\frac{c_{u}c_{0}}{3}b^{k_{i}}\leq|Q_{i}|\leq 4C_{0}b^{k_{i}}.$
So (2.3) and the above ensure that
$\delta^{1/\theta}\leq|Q_{i}|\leq\delta,$
for all $i\in I_{\delta}$. As a result, trivially $\dim_{\theta}E\leq\inf A_{\theta}$.
Fix $s>\dim_{\theta}E$. We will show $s\geq\inf A_{\phi}$ for values $\phi<\theta$ as close to $\theta$ as desired. Then, by letting $s\to\dim_{\theta}E$, and using $\dim_{\phi}E\leq\inf A_{\phi}$, taking $\phi\to\theta$ and using the intermediate dimension’s continuity yields the needed characterization.
For $\varepsilon>0$, there is $\delta_{\varepsilon}=\delta_{\varepsilon}(s)>0$ such that for all $\delta<\delta_{\varepsilon}$ there is a cover $\{U_{i}\}$ of $E$ with
(2.4) $\delta^{1/\theta}\leq|U_{i}|\leq\delta$
and
(2.5) $\sum|U_{i}|^{s}<\varepsilon.$
Let $k_{i}$ be the unique integer such that
(2.6) $4C_{0}b^{k_{i}}\leq|U_{i}|<4C_{0}b^{k_{i}-1}.$
Claim: For every $i\in I_{\delta}$, $U_{i}$ can be covered by at most
$N=C_{d}\Big(\frac{12C_{0}b^{-1}+24C_{0}}{c_{0}}\Big)^{\log_{2}C_{d}}$
cubes of level $k_{i}$.
Proof of Claim.
For the rest of the proof of this Claim, we fix $i\in I_{\delta}$, and set $k=k_{i}$, $U=U_{i}$ to ease the notation.
Set
$J^{U}_{k}=\{j\in J_{k}:Q_{j}^{k}\cap U\neq\emptyset\}$
to be the set of indices of $k$-level cubes that intersect the set $U$. Moreover, set
$Q_{U}=\bigcup_{j\in J^{U}_{k}}Q_{j}^{k},$
to be the largest collection of $k$-level cubes that intersects and covers $U$.
Let $x_{0}\in Q_{U}$. Suppose $x_{0}\in Q_{U}\setminus U$. Then there is $j_{0}\in J^{U}_{k}$ such that
$x_{0}\in Q_{j_{0}}^{k},\quad\text{and}\quad x_{0}^{\prime}\in Q_{j_{0}}^{k}\cap U.$
Let $x\in Q_{U}$. There are two cases to consider:
Case 1: If $x\in U$, then
$d(x_{0},x)$ $\leq d(x_{0},x_{0}^{\prime})+d(x_{0}^{\prime},x)$
$\leq|Q_{j_{0}}^{k}|+|U|$
$\leq 4C_{0}b^{k}+(4C_{0})b^{k-1}$
$=(4C_{0}b^{-1}+4C_{0})b^{k}.$
Case 2: If $x\in Q_{U}\setminus U$, then $\exists j_{x}\in J^{U}_{k}$ such that
$x\in Q_{j_{x}}^{k},\quad x^{\prime}\in Q_{j_{x}}^{k}\cap U.$
Then,
$d(x_{0},x)$ $\leq d(x_{0},x_{0}^{\prime})+d(x_{0}^{\prime},x^{\prime})+d(x^{\prime},x)$
$\leq|Q_{j_{0}}^{k}|+|U|+|Q_{j_{x}}^{k}|$
$\leq 4C_{0}b^{k}+|U|+4C_{0}b^{k}$
$\leq(4C_{0}b^{-1}+8C_{0})b^{k}.$
All the above was under the assumption that the arbitrary point $x_{0}\in Q_{U}$ is not contained in $U$. Similarly, if $x_{0}\in U$, it can be shown that
$d(x_{0},x)\leq(4C_{0}b^{-1}+4C_{0})b^{k},$
for all $x\in Q_{U}$. As a result,
$d(x_{0},x)\leq(4C_{0}b^{-1}+8C_{0})b^{k},$
for all $x_{0},x\in Q_{U}$, implying that
$|Q_{U}|\leq(4C_{0}b^{-1}+8C_{0})b^{k}.$
Thus, there is some $y\in Q_{U}$ such that $Q_{U}\subset B(y,(4C_{0}b^{-1}+8C_{0})b^{k})$. For simplicity, we denote this ball by $B$. Note that every cube in $Q_{U}$, by Theorem A (iii), contains a ball of radius $\tfrac{c_{0}b^{k}}{3}$, and all these balls lie in $B$ and are disjoint. Hence, by the doubling condition of $X$, there are at most
$N=C_{d}^{\frac{(4C_{0}b^{-1}+8C_{0})b^{k}}{c_{0}b^{k}/3}\log_{2}C_{d}}$
such balls of radius $\tfrac{c_{0}b^{k}}{3}$ inside $B$, which implies that there are at most this many $k$-level cubes to cover $U$.
For every set $U_{i}$, there is a collection of $k_{i}$-level cubes $\{Q_{j}^{k_{i}}\}_{j\in J_{i}}$ that cover $U_{i}$. We use the cubes $\mathcal{Q_{\delta}}:=\{Q_{j}^{k_{i}}\}_{\begin{subarray}{c}j\in J_{i}\\ i\in I_{\delta}\end{subarray}}$ to cover $E$.
Note that, by (2.6), we have for all $i\in I_{\delta}$ that
(2.7) $\frac{b\,\delta^{1/\theta}}{4C_{0}}\leq b\,\frac{|U_{i}|}{4C_{0}}\leq b^{k_{i}}\leq\frac{|U_{i}|}{4C_{0}}\leq\frac{\delta}{4C_{0}},$
Because of (2.7), the cover $\mathcal{Q_{\delta}}$ might not be admissible for the conditions of $A_{\theta}$. Thus, we need to choose a $\phi$ for which $b^{k_{i}}$ satisfies the appropriate lower bound for $\mathcal{Q_{\delta}}$ to be an admissible collection for the conditions in $A_{\phi}$. Set
(2.8) $\phi:=\frac{\theta\log\delta_{\varepsilon}}{\log\delta_{\varepsilon}+\log\frac{b\,c_{u}\,c_{0}}{12C_{0}}},$
and note that $\phi<\theta$. With this choice we have
$\delta_{0}^{\frac{1}{\phi}-\frac{1}{\theta}}=\frac{bc_{u}c_{0}}{12C_{0}},$
which ensures that for all $\delta\leq\delta_{0}$ we have
$\frac{3}{c_{u}c_{0}}\delta^{1/\phi}\leq\frac{b}{4C_{0}}\delta^{1/\theta}.$
Thus, (2.7), (2.8) imply
$\frac{3}{c_{u}c_{0}}\delta^{1/\phi}\leq b^{k_{i}}\leq\frac{\delta}{4C_{0}}$
$b^{k_{i}}\geq$, which makes the collection $\{Q_{i}^{k_{i}}\}$ admissible for $\inf A_{\phi}$.
Also, due to the Claim,
$\sum_{i,j}|Q_{j}^{k_{i}}|^{s}\leq N\;\sum_{i}|U_{i}|^{s}<N\;\varepsilon:=\varepsilon^{\prime}.$
As a result, because $\phi\to\theta$ as $\delta_{\varepsilon}\to 0$, we have shown that for every tiny $\lambda>0$, for every $\varepsilon^{\prime}>0$, there is $\delta_{\varepsilon^{\prime}}^{\prime}$ small enough (as dictated by (2.8) and the value $\delta_{N^{-1}\varepsilon^{\prime}}$) such that for all $\delta\in(0,\delta^{\prime}_{\varepsilon^{\prime}})$, there is a cover $\{Q_{i}\}_{i\in I_{\delta}}$ of $E$ by cubes of level $k_{i}$ with
$\frac{3}{c_{u}c_{0}}\delta^{1/(\theta+\lambda)}\leq b^{k_{i}}\leq\frac{1}{4C_{0}}\delta,$
so that
$\sum_{i}|Q_{i}|^{s}<\varepsilon.$
This implies
$\dim_{\theta+\lambda}E\leq\inf A_{\theta+\lambda}\leq s.$
Since $s>\dim_{\theta}E$ was arbitrary, we let $s\rightarrow\dim_{\theta}E$, and then letting $\lambda\to 0$, by continuity of intermediate dimensions, completes the proof. ∎
Remark 2.3.
The construction of dyadic cubes in [52] was actually given for quasimetric spaces. As a result, Proposition 2.2 is also true if $X$ is a quasimetric doubling space. The proof is almost identical, with the only difference being the dependence of a few of the constants on the quasimetric constant of the space.
2.2. Mappings between metric spaces
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two metric spaces. Given $\alpha>0$, a mapping $f:X\rightarrow Y$ and a set $B\subset X$, we define the $\alpha$-Hölder coefficient of $f$ on $B$ as
$|f|_{\alpha,B}:=\sup\left\{\frac{d_{Y}(f(x),f(y))}{[d_{X}(x,y)]^{\alpha}}:\,x,y\in B\,\,\text{distinct}\right\}.$
If $|f|_{\alpha,B}<\infty$ then we say that $f$ is $\alpha$-Hölder continuous in $B$.
Given an at most countable index set $I$, we denote by $\ell^{p}(I)$ the space of real-valued sequences $\{c_{i}\}_{i\in I}$ with finite $p$-norm $(\sum_{i\in I}c_{i}^{p})^{1/p}<\infty$. We call $\sum_{i\in I}c_{i}^{p}$ the $p$-sum of the sequence $\{c_{i}\}_{i\in\mathbb{N}}$.
For the rest of the paper, all index sets are assumed to be at most countable. We now recall the class of compactly Hölder mappings.
Definition 2.4.
Let $f:X\rightarrow Y$ be a mapping between two arbitrary metric spaces. For $p\in(1,\infty)$ and $\alpha>0$, we say $f$ is $(p,\alpha)$-compactly Hölder, and write $f\in CH^{p,\alpha}(X:Y)$, if for any compact set $E\subset X$ and any $\varepsilon\in(0,1)$ there are $r_{E}>0$ and $C_{E}>0$ satisfying the following:
if $\{B_{i}\}_{i\in I}$ is a collection of balls $B_{i}:=B(x_{i},r)$ with $x_{i}\in X$, $r<r_{E}$ that covers $E$ and $B(x_{i},\varepsilon r)\cap B(x_{j},\varepsilon r)=\emptyset$ for all distinct $i,j\in I$, then the $p$-sum of the Hölder coefficients of $f$ on $B_{i}$ is at most $C_{E}$, i.e.,
(2.9) $\sum\limits_{i\in I}|f|_{\alpha,B_{i}}^{p}\leq C_{E}.$
Here we follow the convention that if $\{B(x_{i},r)\}_{i\in I}$ covers $E$, it is implied that $B(x_{i},r)\cap E\neq\emptyset$ for all $i$, but not all $x_{i}$ necessarily lie in $E$. Note that applying the definition on singleton sets yields that compactly Hölder mappings are Hölder continuous on compact sets. Moreover, in the setting of Definition 2.4 it is actually implied by (2.9) that there are $C_{i}>0$ such that
(2.10) $|f(B(x_{i},r)|\leq C_{i}|B(x_{i},r)|^{\alpha}$
with $\sum_{i\in I}C_{i}^{p}\leq C_{E}$. This inequality provides insight on the relation with the Euclidean setting. More specifically, it hints at how the motivation for Definition 2.4 comes from continuous super-critical Sobolev maps between Euclidean spaces, i.e. continuous maps in $W^{1,p}(\Omega;\mathbb{R}^{n})$ with $\Omega\subset\mathbb{R}^{n}$ and $p>n$. For more details on the motivation behind these mappings see [21].
Remark 2.5.
Given $f\in CH^{p,\alpha}(X:Y)$, a non-empty compact set $E\subset X$, some $\varepsilon\in(0,1)$ and a cover of $E$ by balls $\{B(x_{i},r_{i})\}_{i\in I}$ with $r_{i}\leq r_{E}$ and $B(x_{i},\varepsilon r_{i})\cap B(x_{j},\varepsilon r_{j})=\emptyset$ for all distinct $i,j\in I$, then there are $C_{i}>0$ with
(2.11) $|f(B(x_{i},r_{i})|\leq C_{i}|B(x_{i},r_{i})|^{\alpha}$
and $\sum_{i\in I}C_{i}^{p}\leq C_{E}$. This change from radii $r$ to not necessarily equal radii $r_{i}$ can be made by bounding the partial sum $\sum_{i\in\{i_{1},\dots i_{n}\}}|f|_{\alpha,B(x_{i},r_{i})}$ by a sum of the $\alpha$-Hölder semi-norms over balls $B(y_{j},\rho)$, $\rho=\min\{r_{i}:i\in\{i_{1},\dots i_{n}\}\}$, which cover $\cup_{i\in\{i_{1},\dots i_{n}\}}B(x_{i},r_{i})$ and are $\varepsilon$-disjoint, with the latter sum being by definition at most $C_{E}$, for all $n\in\mathbb{N}$.
[…]
Let $(X,d_{X},\mu)$ be a metric measure space, $(Y,d_{Y})$ be a metric space equipped with the Borel sigma-algebra, and $s\in(0,\infty)$. Following [38, 39, 64], a measurable function $g:X\to[0,\infty]$ is called an $s$-gradient of a measurable function $u\colon X\rightarrow Y$ if there exists a set $E\subset X$ with $\mu(E)=0$ such that
(2.12) $d_{Y}(u(x),u(y))\leq[d_{X}(x,y)]^{s}\left[g(x)+g(y)\right],$
for every $x,y\in X\setminus E$. The collection of all the $s$-gradients of $u$ is denoted by $\mathcal{D}^{s}(u)$. Given $p\in(0,\infty)$, the (homogeneous) fractional Hajłasz–Sobolev space $\dot{M}^{s,p}(X:Y)$ is defined as the collection of all the measurable functions $u\colon X\rightarrow Y$ such that
$\|u\|_{\dot{M}^{s,p}(X:Y)}:=\inf_{g\in\mathcal{D}^{s}(u)}\|g\|_{L^{p}(X)}<\infty.$
Here and thereafter, we make the agreement that $\inf\emptyset:=\infty$.
In order to define the Hajłasz–Triebel–Lizorkin and Hajłasz–Besov spaces, we need a suitable notion for the gradient. Following [57], a sequence $\{g_{k}\}_{k\in\mathbb{Z}}$ of measurable functions $g_{k}:X\to[0,\infty]$ is called a fractional $s$-gradient of a measurable function $u\colon X\rightarrow Y$ if there exists a set $E\subset X$ with $\mu(E)=0$ such that
(2.13) $d_{Y}(u(x),u(y))\leq[d_{X}(x,y)]^{s}\left[g_{k}(x)+g_{k}(y)\right]$
for any $k\in\mathbb{Z}$ and $x,y\in X\setminus E$ satisfying $2^{-k-1}\leq d_{X}(x,y)<2^{-k}.$ Let $\mathbb{D}^{s}(u)$ denote the set of all the fractional $s$-gradients of $u$. Given $p\in(0,\infty)$, $q\in(0,\infty]$, and a sequence $\vec{g}:=\{g_{k}\}_{k\in\mathbb{Z}}$ of measurable functions $g_{k}:X\to[0,\infty]$, define
$\|\vec{g}\|_{L^{p}(X,\ell^{q})}:=\left\|\,\|\{g_{k}\}_{k\in\mathbb{Z}}\|_{\ell^{q}}\right\|_{L^{p}(X)}$
and
$\|\vec{g}\|_{\ell^{q}(L^{p}(X))}:=\left\|\left\{\|g_{k}\|_{L^{p}(X)}\right\}_{k\in\mathbb{Z}}\right\|_{\ell^{q}},$
where
$\|\{g_{k}\}_{k\in\mathbb{Z}}\|_{\ell^{q}}:=\begin{cases}\displaystyle\left(\sum_{k\in\mathbb{Z}}|g_{k}|^{q}\right)^{1/q}&~\text{if}~q\in(0,\infty),\\ \displaystyle\sup_{k\in\mathbb{Z}}|g_{k}|&~\text{if}~q=\infty.\end{cases}$
Then the (homogeneous) Hajłasz–Triebel–Lizorkin space $\dot{M}^{s}_{p,q}(X:Y)$ is defined as the collection of all the measurable mappings $u\colon X\rightarrow Y$ such that the semi-norm
$\|u\|_{\dot{M}^{s}_{p,q}(X:Y)}:=\inf_{\vec{g}\in\mathbb{D}^{s}(u)}\|\vec{g}\|_{L^{p}(X,\ell^{q})}<\infty.$
The (homogeneous) Hajłasz–Besov space $\dot{N}^{s}_{p,q}(X:Y)$ is defined as the collection of all the measurable mappings $u\colon X\rightarrow Y$ such that the semi-norm
$\|u\|_{\dot{N}^{s}_{p,q}(X:Y)}:=\inf_{\vec{g}\in\mathbb{D}^{s}(u)}\|\vec{g}\|_{\ell^{q}(L^{p}(X))}<\infty.$
A few comments are in order. We use the term ‘semi-norms’ for $\|\cdot\|_{\dot{M}^{s}_{p,q}(X:Y)}$, $\|\cdot\|_{\dot{M}^{s}_{p,q}(X:Y)}$, and $\|\cdot\|_{\dot{M}^{s}_{p,q}(X:Y)}$ even though the triangle inequality holds only when $p,q\geq 1$. A genuine ‘norm’ can be obtained by passing to the quotient space modulo constant functions. Since altering functions on sets of measure zero do not affect their membership in $\dot{M}^{s,p}(X:Y)$, $\dot{M}^{s}_{p,q}(X:Y)$, or $\dot{N}^{s}_{p,q}(X:Y)$, it is standard to regard these spaces as consisting of equivalence classes of functions. We adopt this convention here, but we choose to omit the details.
When $Y=\mathbb{R}$, we make the following abbreviations: $\dot{M}^{s,p}(X):=\dot{M}^{s,p}(X:\mathbb{R})$ $\dot{M}^{s}_{p,q}(X):=\dot{M}^{s}_{p,q}(X:\mathbb{R})$, and $\dot{N}^{s}_{p,q}(X):=\dot{N}^{s}_{p,q}(X:\mathbb{R})$. It was shown in [57] that $\dot{M}^{s}_{p,q}(\mathbb{R}^{n})$ coincides with the classical Triebel–Lizorkin space $\dot{F}^{s}_{p,q}(\mathbb{R}^{n})$ for any $s\in(0,1)$, $p\in(\frac{n}{n+s},\infty)$, and $q\in(\frac{n}{n+s},\infty]$, and $\dot{N}^{s}_{p,q}(\mathbb{R}^{n})$ coincides with the classical Besov space $\dot{B}^{s}_{p,q}(\mathbb{R}^{n})$ for any $s\in(0,1)$, $p\in(\frac{n}{n+s},\infty)$, and $q\in(0,\infty]$. In particular, the Hajłasz–Triebel–Lizorkin and Hajłasz–Besov spaces on $\mathbb{R}^{n}$ contain the classical fractional Sobolev spaces as special cases (see [37, Chapter 2]). When $s=1$, we have that $\dot{M}^{1}_{p,\infty}(\mathbb{R}^{n})=\dot{M}^{1,p}(\mathbb{R}^{n})$ coincides with the classical Sobolev space $\dot{W}^{1,p}(\mathbb{R}^{n})$ for any $p\in(1,\infty)$; see Lemma 4.5 and [39]. We refer the reader to [38, 39, 64, 57, 43, 3, 42, 5, 4] for more information on Sobolev, Triebel–Lizorkin, and Besov spaces on metric measure spaces.
The following notions for measures are also typically needed in this setting. We say a metric measure space $(X,d,\mu)$ is locally $Q$-homogeneous, for some $Q>0$, if for all compact $K\subset X$ there are constants $\tilde{R}_{\text{hom}}(K)>0$, $\tilde{C}_{\text{hom}}(K)\geq 1$ such that
(2.14) $\frac{\mu(B(x,r_{2}))}{\mu(B(x,r_{1}))}\leq\tilde{C}_{\text{hom}}(K)\left(\frac{r_{2}}{r_{1}}\right)^{Q},$
for all $x\in K$ and scales $0<r_{1}<r_{2}<\tilde{R}_{\text{hom}}(K)$. Note that local $Q$-homogeneity implies that $(X,d,\mu)$ is locally doubling (see[12]), i.e., for every compact subset $K\subseteq X$, there exists a radius $R>0$ and a constant $C\geq 1$ such that
$\mu(B(x,2r))\leq C\,\mu(B(x,r))$
whenever $B(x,r)$ is a ball centered at a point in $K$ with $r\leq R$. We say a metric space $(X,d)$ is locally $Q$-homogeneous if there is a measure $\mu$ on $X$ such that $(X,d,\mu)$ is locally $Q$-homogeneous. One particular property due to local homogeneity that we need is the lower bound on the measure of a ball by its radius to a power. More specifically, for $R_{\text{hom}}(K)=\tilde{R}_{\text{hom}}(K)/3$ and a potentially larger $C_{\text{hom}}(K)$, it can be shown that
(2.15) $\frac{r^{Q}}{C_{\text{hom}}(K)}\leq\mu(B(x,r)),$
for all $x\in K$ and $r\in(0,R_{\text{hom}}(K))$.
It can be necessary at times to also have a similar upper bound on the measure. We say that $(X,d,\mu)$ is $Q$-Ahlfors regular for some $Q>0$ if there is a constant $C_{A}>0$ such that for all $x\in X$ and all $r\in(0,|X|)$ we have
$\frac{1}{C_{A}}r^{Q}\leq\mu(B(x,r))\leq C_{A}r^{Q}.$
We say a metric space $(X,d)$ is $Q$-Ahlfors regular if there is a measure $\mu$ on $X$ such that $(X,d,\mu)$ is $Q$-Ahlfors regular. Note that $Q$-regularity of a measure implies the $Q$-homogeneous property.
3. Intermediate dimensions under compactly Hölder mappings
Suppose $(Y,d_{Y})$ is a $c_{u}^{\prime}$-uniformly perfect metric space and $(X,d_{X})$ is a $c_{u}$-uniformly perfect, $C_{d}$-doubling metric space with a fixed system of dyadic cubes as in Section 2. We further assume that both $X$ and $Y$ have more than one point each, as otherwise all results trivially hold. Before we proceed, we need the following elementary property of uniformly perfect spaces.
Lemma 3.1.
If $Y$ is uniformly perfect with constant $c_{u}^{\prime}>0$, then for every $F\subset Y$ and every $M>0$ with $|F|<M<Mc_{u}^{\prime-1}<|Y|$, there is a set $L(F)$ containing $F$ with
$M\leq|L(F)|\leq 2c_{u}^{\prime-1}M.$
Proof.
Let $y\in F$, then $B_{Y}(y,Mc_{u}^{\prime-1})$ contains $F$, and
$|B_{Y}(y,Mc_{u}^{\prime-1})|\geq c_{u}^{\prime}c_{u}^{\prime-1}M=M$
and
$|B_{Y}(y,Mc_{u}^{\prime-1})|\leq 2Mc_{u}^{\prime-1}.$
Hence, $L(F)=B_{Y}(y,Mc_{u}^{\prime-1})$ is the desired set. ∎
Since all mappings considered henceforth are defined on $X$ and into $Y$, we set $CH^{p,\alpha}=CH^{p,\alpha}(X:Y)$. Let $f\in CH^{p,\alpha}$ for $p\in(1,\infty)$, $\alpha>0$ and $E\subset X$ be a bounded set. We plan on using coverings of $E$ by dyadic cubes, provided by Proposition 2.2, and determine admissible coverings for $\dim_{\theta}f(E)$ based on those of $E$ and shrinking properties of $f$ controlled by inequality (2.11). Before we delve into the proof, we give an outline of the combinatorial part of our method. The main idea is to fix an arbitrary $d>\dim_{\theta}E$, and for every $\varepsilon>0$ fix for every $\delta\in(0,\delta_{\varepsilon})$ a dyadic covering $\mathcal{Q}_{\delta}$ of $E$ as in Proposition 2.2. The images of the covering cubes form a cover of $f(E)$. After determining a $\delta_{Y}$ to use for $\dim_{\theta}f(E)$ (see (3.3)), we need to modify the cubes in order to get a $\delta_{Y}^{1/\theta}$-admissible cover of $f(E)$. To do so, we build a combinatorial graph with cubes as vertices, in the following manner (see Figure 1):
  • Let $\{m_{1},\dots,m_{N_{\delta}}\}$ be the set of all levels of cubes in $\mathcal{Q}_{\delta}$, with $m_{i}<m_{j}$ for all $i<j$. List all $m_{1}$-level cubes in $\mathcal{Q}_{\delta}$ on the first row, all $m_{2}$-level cubes in $\mathcal{Q}_{\delta}$ $m_{2}-m_{1}$ rows lower, all $m_{3}$-level cubes in $\mathcal{Q}_{\delta}$ $m_{3}-m_{2}$ rows lower and so on, until the highest level $m_{N_{\delta}}$ for cubes in $\mathcal{Q}_{\delta}$.
  • Color all cubes whose image under $f$ has diameter larger than $\delta_{Y}$ red, those with images of diameter less than $\delta_{Y}^{1/\theta}$ green, and the remaining blue.
  • Subdivide every red cube, based on Theorem A, into next level sub-cubes, and draw in the next row from their ancestor those that intersect $E$, and connect them with edges to their ancestor. These descendant cubes could be red, blue, or green, depending on the behavior of $f$.
  • Iterate the previous step, until we reach the first level $m_{K_{\delta}}\geq m_{N_{\delta}}$ with no red cubes. This is guaranteed by uniform continuity of $f$.
Figure 1. The graph $\mathcal{G}(\mathcal{Q}_{\delta})$ above includes a few noteworthy characteristics of $\mathcal{Q}_{\delta}$ and $f$. It is depicted that $\mathcal{Q}_{\delta}$ has no cubes of level $m_{1}+2$, and while it has no cubes of level $m_{1}+4$ either, we need to subdivide red cubes to and past that level for the induced desired covering of $f(E)$. Moreover, there are red cubes of $\mathcal{Q}_{\delta}$ even at the very last level, which means that $f$ is capable of increasing the diameter of even these small cubes, creating the need to subdivide beyond the $m_{N_{\delta}}$ level.
Based on the above graph, the strategy then is to use a covering of $f(E)$ consisting of images of all blue cubes, and “enlarged” images of green cubes (using Lemma 3.1). This by construction is a $\delta_{Y}^{1/\theta}$-admissible covering of $f(E)$, with the least cardinality possible based on $\mathcal{Q}_{\delta}$, and with a combinatorial representation that facilitates estimates on the sum of respective diameters to an appropriate exponent, achieving the desired upper bound on $\dim_{\theta}f(E)$. With this strategy in mind, we are ready to delve into the proof of Theorem 1.1.
Proof of Theorem 1.1.
We focus on the case $d_{E}(\theta)\leq pd_{E}(\theta)(\alpha p+d_{E}(\theta))^{-1}$, and the other case can be treated similarly. Suppose $\dim_{\theta}E=d_{E}<d<d^{\prime}<p-\alpha p$, and set $D=\tfrac{p{d^{\prime}}}{\alpha p+d^{\prime}}$. The choice $d<d^{\prime}<p-\alpha p$ guarantees that $d\neq D$, but the proof is similar in the case $d_{E}\geq p-\alpha p$, by choosing $d>d_{E}$.
Due to $d>\dim_{\theta}E$ and Proposition 2.2, for $\varepsilon>0$ there is $\delta_{\varepsilon}\in(0,1)$, such that for all $\delta\in(0,\delta_{\varepsilon})$ there is a covering $\mathcal{Q_{\delta}}:=\{Q_{i}^{k_{i}}\}_{i\in I_{\delta}}$ of $E$ by dyadic cubes such that
(3.1) $\frac{3}{c_{u}c_{0}}\,\delta^{1/\theta}\leq b^{k_{i}}\leq\frac{1}{4C_{0}}\,\delta,$
for all $i\in I_{\delta}$, and
(3.2) $\sum_{i\in I_{\delta}}|Q_{i}^{k_{i}}|^{d}<\varepsilon.$
Without loss of generality, we may assume that $\delta_{\varepsilon}$ is small enough for certain properties to apply. Namely, we assume that
  • $\delta_{\varepsilon}<\varepsilon<1$, in order to replace $\delta$ by $\varepsilon$ when all that matters is for a quantity to be small,
  • $\delta_{\varepsilon}<r_{E}$, to ensure that (2.11) can be applied to the corresponding balls $B(Q_{i}^{k_{i}})$ from Theorem A (iii) for the cubes in $\mathcal{Q}_{\delta}$, and all their sub-cubes,
  • $\delta_{\varepsilon}<2^{-1}\min\{|X|,(C_{E}^{-1}|Y|)^{1/\alpha}\}$, to ensure that the uniformly perfect properties of $X$ and $Y$ can be applied to balls $B(Q_{i}^{k_{i}})$ in $X$, and balls of radius $|f(Q_{i}^{k_{i}})|$ in $Y$, respectively.
  • $\delta_{\varepsilon}<(2^{-1}|Y|/c_{u}^{\prime})^{D/d}$, to ensure that the uniformly perfect property of $Y$ can be applied to balls of radius $\delta^{d/(D\theta)}/c_{u}^{\prime}$ in $Y$ (i.e., Lemma 3.1 applies for sets $F$ with $|F|<M=\delta^{d/(D\theta)}$).
Without loss of generality also assume $k_{1}\leq k_{2}\leq\;s\leq k_{N_{\delta}}$. Moreover, if $Q_{i^{\prime}}^{k_{i^{\prime}}}\subset Q_{i}^{k_{i}}$ for $i^{\prime}\neq i\in I_{\delta}$, then we can reduce $\{Q_{i}^{k_{i}}\}_{i\in I_{\delta}}$ to $\{Q_{i}^{k_{i}}\}_{i\in I_{\delta}\setminus\{i^{\prime}\}}$, and (3.1), (3.2) would still be true for $i\in I_{\delta}\setminus\{i^{\prime}\}$, since $Q_{i^{\prime}}^{k_{i^{\prime}}}$ would be redundant for the covering. As a result, we may assume that if $Q_{i}^{k_{i}}$ is a cube in the cover $\mathcal{Q_{\delta}}$, then no sub-cube of $Q_{i}^{k_{i}}$ is contained in $\mathcal{Q_{\delta}}$. Note that $k_{i}$’s are not necessarily pairwise distinct, since, for instance, there could be two level $m=k_{1}=k_{2}$ cubes in $\mathcal{Q_{\delta}}$. We consider a re-labeling of the cubes in $\mathcal{Q_{\delta}}$ to account for that, namely, set
$I_{\delta}^{\prime}:=\{k_{i}:i\in I_{\delta}\},$
which can be represented as $I_{\delta}^{\prime}=\{m_{1},m_{2},\dots,m_{N_{\delta}}\}$, with positive integers $m_{i}$, with $m_{i}<m_{j}$ for all $i<j$. This notation helps by clarifying exactly which level we address in following arguments. We also set
(3.3) $\delta_{Y}:=\delta^{d/D}.$
We say a cube $Q$ that intersects $E$ and is contained in some $Q_{i}^{k_{i}}\in\mathcal{Q}_{\delta}$, for some $i\in I_{\delta}$, is blue if $|f(Q)|\in[\delta_{Y}^{1/\theta},\delta_{Y})$, green if $|f(Q)|<\delta_{Y}^{1/\theta}$, and red otherwise. Notice that the terminology applies not only to cubes in $\mathcal{Q_{\delta}}$, but also to their sub-cubes. In addition, it applies only to the aforementioned types of cubes, due to our initial reduction on $\mathcal{Q}_{\delta}$ and by Theorem A (i). By uniform continuity of $f$, it is guaranteed that past a certain level, all sub-cubes of cubes in $\mathcal{Q_{\delta}}$ will be blue or green. The strategy is to sub-divide every red cube in $\mathcal{Q_{\delta}}$ and all their red sub-cubes the least number of times necessary, in order for the resulting cubes to all be blue or green. Then, apply Lemma 3.1 to the images of all green cubes under $f$, in order to “enlarge” them, and pick a covering of $f(E)$ consisting of these enlarged sets and the images of all blue cubes under $f$. By construction, this covering is a $\delta_{Y}^{1/\theta}$-admissible covering for $\dim_{\theta}f(E)$. Then, it would be enough to show that the sum of their diameters is small enough.
Suppose $m_{K_{\delta}}\geq m_{N_{\delta}}$ is the smallest integer such that all cubes of level $m_{K_{\delta}}$ which intersect $E$ and are contained in $\bigcup_{i\in I_{\delta}}Q_{i}^{k_{i}}$, are either green or blue. Note that $m_{K_{\delta}}$ could be strictly larger than $m_{N_{\delta}}$, because we can have red $m_{N_{\delta}}$-level cubes in $\mathcal{Q_{\delta}}$, or red cubes in $\mathcal{Q_{\delta}}$ might need to be subdivided past the level $m_{N_{\delta}}$ to give only blue and green descendants. Based on this subdivision of red cubes in $\mathcal{Q_{\delta}}$ and their red sub-cubes, solely dependent on the way $f$ distorts diameters, we build a directed graph $\mathcal{G}=\mathcal{G}(\mathcal{Q_{\delta}})$. For an integer $m\in[m_{1},m_{K_{\delta}}]$, set $V_{m}$ to be the collection of all $m$-level cubes intersecting $E$ and contained in $\cup_{i\in I_{\delta}}Q_{i}^{k_{i}}$ that are either red, or are contained in a red cube of level m1m-1. Then, the vertex set of $\mathcal{G}$ is defined to be the collection of cubes
$V=\mathcal{Q}_{\delta}\cup\bigcup_{m=m_{1}}^{m_{K_{\delta}}}V_{m},$
and the edges of $\mathcal{G}$ are
$E=\bigcup_{m=m_{1}}^{m_{K_{\delta}}-1}\big\{(Q_{t},Q_{s}):\;Q_{t}\in\mathcal{D}_{m},Q_{s}\in\mathcal{D}_{m+1},\;Q_{s}\subsetneq Q_{t},\;\text{and }Q_{t}\text{ red}\big\}.$
What we use in the desired covering of $f(E)$ is essentially the images of blue vertices in $V$ and the “enlarged” images of green vertices in $V$ (after applying Lemma 3.1). For an integer $m\in[m_{1},m_{K_{\delta}}]$, we call the collection of all $m$-level cubes in $V$, i.e. the collection $\mathcal{Q}_{\delta}\cup V_{m}$, the $m$-th row in $\mathcal{G}$. (See also Figure 1).
Note that blue and green cubes in $\mathcal{Q_{\delta}}$ are part of no edge, because there is no need for them to be sub-divided, and they cannot have an ancestor also in $\mathcal{Q}_{\delta}$ by the respective reduction in the beginning of the proof. Let $V_{R},V_{B}^{\text{orig}},V_{B}^{\prime},V_{G}$ denote the red cubes in $V$, blue cubes in $V\cap\mathcal{Q_{\delta}}$, blue cubes in $V$ that do not lie in $\mathcal{Q_{\delta}}$ (i.e. they have an ancestor in $V$), and green cubes in $V$, respectively. Notice that if we have a bound on $\text{card}V_{R}$, we can also bound $\text{card}V_{B}^{\prime}$ by Remark 2.1:
(3.4) $\text{card}V_{B}^{\prime}\leq N_{d}\;\text{card}V_{R},$
because every vertex in $V_{B}^{\prime}$ is connected to a red vertex in $V$, and for each red $Q\in V$ there are at most $N_{d}$ cubes $\tilde{Q}\in V$ with $(Q,\tilde{Q})\in E$.
Note, by (3.1) and (3.2), that
$\text{card}\mathcal{Q_{\delta}}\leq c_{u,0}\,\varepsilon\,\delta^{-d/\theta},$
where $c_{u,0}=(\frac{c_{u}c_{0}}{3})^{d}$. This implies
(3.5) $\text{card}V_{G}\leq 2\max\Big\{c_{u,0}\,\varepsilon\,\delta^{-d/\theta},\;N_{d}\;\text{card}V_{R}\Big\},$
where $c_{u,0}\varepsilon\delta^{-d/\theta}$ bounds the number of green cubes in $\mathcal{Q}_{\delta}$, and $N_{d}\,\text{card}V_{R}$ bounds the number of green cubes in $V$ which are connected to some red vertex in $V$ by an edge, after another application of Remark 2.1. Let
$\mathcal{U}=\mathcal{U}_{\delta}=\{f(Q):Q\in V_{B}^{\text{orig}}\cup V_{B}^{\prime}\}\cup\{L(f(Q)):Q\in V_{G}\},$
where $L(f(Q))$ is a set as in Lemma 3.1 applied to $f(Q)$ for $M=\delta_{Y}^{1/\theta}$. By definition of blue and green cubes and by Lemma 3.1, which is applicable due to assuming that $\delta_{\varepsilon}$ is small enough so that $2c_{u}^{\prime-1}\delta_{Y}^{1/\theta}\leq\delta_{Y}$ for $\delta\leq\delta_{\varepsilon}$ (see last point after (3.2)), the collection $\mathcal{U}$ is a $\delta_{Y}^{1/\theta}$-admissible cover of $f(E)$. Thus, we need to estimate the sum
$\mathcal{S}:=\sum_{U\in\mathcal{U}}|U|^{D}=\sum_{Q\in V_{B}^{\text{orig}}}|f(Q)|^{D}+\sum_{Q\in V_{B}^{\prime}}|f(Q)|^{D}+\sum_{Q\in V_{G}}|L(f(Q))|^{D}.$
But, by definition of blue cubes and (3.4),
$\sum_{Q\in V_{B}^{\prime}}|f(Q)|^{D}\leq\sum_{Q\in V_{B}^{\prime}}\delta_{Y}^{D}\leq N_{d}\;\text{card}(V_{R})\;\delta_{Y}^{D},$
and
$\sum_{Q\in V_{G}}|L(f(Q))|^{D}\leq 2\max\Big\{c_{u,0}\,\varepsilon\,\delta^{-d/\theta},\;N_{d}\;\text{card}V_{R}\Big\}\,\delta_{Y}^{D/\theta},$
which by the above imply
(3.6) $\mathcal{S}\leq\sum_{Q\in V_{B}^{\text{orig}}}|f(Q)|^{D}+N_{d}\text{card}(V_{R})\,\delta_{Y}^{D}+2\max\Big\{c_{u,0}\,\varepsilon\,\delta^{-d/\theta},\;N_{d}\text{card}V_{R}\Big\}\delta_{Y}^{D/\theta}.$
We estimate $\sum_{Q\in V_{B}^{\text{orig}}}|f(Q)|^{D}$ using $f\in CH^{p,\alpha}$, the defining properties of compactly–Hölder maps, and Hölder’s inequality, noting that $p/D=\alpha p/d+1>1$, and
$\frac{1}{p/D}+\frac{1}{\tfrac{p}{p-D}}=1.$
Namely,
$\sum_{Q\in V_{B}^{\text{orig}}}|f(Q)|^{D}$ $\leq\sum_{Q\in V_{B}^{\text{orig}}}|f(B(Q))|^{D}$
$\leq\sum_{Q\in V_{B}^{\text{orig}}}C_{Q}^{D}|B(Q)|^{D\alpha}$
$\lesssim\sum_{Q\in V_{B}^{\text{orig}}}C_{Q}^{D}|Q|^{D\alpha}$
(3.7) $\leq\left(\sum_{Q\in V_{B}^{\text{orig}}}C_{Q}^{p}\right)^{D/p}\left(\sum_{Q\in V_{B}^{\text{orig}}}|Q|^{D\alpha\tfrac{p}{p-D}}\right)^{\tfrac{p-D}{p}}.$
The comparability constant above (3.7) depends only on the fixed constants $p,\alpha,d^{\prime}$ and the uniform constants $c_{0},C_{0},b$ from Theorem A. Note that
$D\alpha\frac{p}{p-D}=\frac{\alpha p^{2}d^{\prime}(\alpha p+d^{\prime})^{-1}}{p-(\alpha p+d^{\prime})^{-1}pd^{\prime}}=d^{\prime},$
and
$\sum_{Q\in V_{B}^{\text{orig}}}|Q|^{d^{\prime}}=\sum_{Q\in V_{B}^{\text{orig}}}\Big(\tfrac{|Q|}{\delta}\Big)^{d^{\prime}}\delta^{d^{\prime}}\leq\delta^{d^{\prime}}\sum_{Q\in V_{B}^{\text{orig}}}\Big(\tfrac{|Q|}{\delta}\Big)^{d},$
due to (3.1) and $d^{\prime}>d$. Thus, (3.7) yields
$\sum_{Q\in V_{B}^{\text{orig}}}|f(Q)|^{D}\leq(C_{E})^{D/p}\;\left(\delta^{\,d^{\prime}-d}\,\sum_{Q\in V_{B}^{\text{orig}}}|Q|^{d}\right)^{(p-D)/p}\lesssim\delta^{\tfrac{(d^{\prime}-d)(p-D)}{p}}\,\varepsilon^{\tfrac{p-D}{p}},$
where the comparability constant similarly only depends on uniformly fixed constants. Since $d^{\prime}>d$ and $p-D=\alpha p^{2}(\alpha p+d^{\prime})^{-1}>0$, and $\delta<1$, the above gives
(3.8) $\sum_{Q\in V_{B}^{\text{orig}}}|f(Q)|^{D}\leq\varepsilon^{(p-D)/p}.$
For the remaining terms in (3.6), it is enough to bound the number of red vertices of $\mathcal{G}$. Let $M(m)$ denote the number of red vertices in the $m$-th row of $\mathcal{G}$.
Let $\{Q_{j}^{m}\}_{j\in J_{m}}$ be all red $m$-level cubes. If $Q_{j}^{m}$ is such a cube, by $f\in CH^{p,\alpha}$ and (2.11) we have
$|f(Q_{j}^{m})|\leq|f(B(Q_{j}^{m}))|\leq C_{Q_{j}^{m}}\,|B(Q_{j}^{m})|^{\alpha}.$
But by $|f(Q_{j}^{m})|\geq\delta_{Y}$, and $|B(Q_{j}^{m})|\simeq|Q_{j}^{m}|\simeq b^{m}$, due to Theorem A (iii) and uniform perfectness of $X$, we get
(3.9) $\delta_{Y}^{p}\lesssim C_{Q_{j}^{m}}^{p}\;b^{m\alpha p}.$
Applying this to all level $m$ red vertices and summing (3.9) over all of them gives
$M(m)\lesssim\delta_{Y}^{-p}\sum_{j\in J_{m}}C_{Q_{j}^{m}}^{p}\,b^{m\alpha p}\leq C_{E}\,b^{m\alpha p}\,\delta_{Y}^{-p}.$
Summing the above over all levels $m\geq m_{1}$ yields
$\text{card}V_{R}=\sum_{m=m_{1}}^{m_{K}}M(m)\lesssim C_{E}\,\delta_{Y}^{-p}\,\sum_{m\geq m_{1}}b^{m\alpha p}\lesssim\delta_{Y}^{-p}\,b^{m_{1}\alpha p}.$
But by $k_{1}=m_{1}$ and (3.1) we have
$\text{card}V_{R}\lesssim\delta_{Y}^{-p}\,\delta^{\alpha p}.$
Using the above and (3.8) on (3.6) yields
$\mathcal{S}\lesssim\varepsilon^{(p-D)/p}+N_{d}\,\delta_{Y}^{-p}\,\delta^{\alpha p}\,\delta_{Y}^{D}+2\max\left\{c_{u,0}\,\varepsilon\,\delta^{-d/\theta},\;N_{d}\,\delta_{Y}^{-p}\,\delta^{\alpha p}\right\}\,\delta_{Y}^{D/\theta}.$
By choice of $\delta_{Y}$ in (3.3), $\delta_{Y}=\delta^{d/D}$, we have
$N_{d}\,\delta_{Y}^{-p}\,\delta^{\alpha p}\,\delta_{Y}^{D}=N_{d}\,\delta^{-dp/D}\,\delta^{\alpha p}\,\delta^{d},$
and because $dp/D+\alpha p+d=(1-d/d^{\prime})\alpha p>0$, and
$dp/D+\alpha p+d/\theta=(1-d/d^{\prime})\alpha p+(1/\theta-1)d>0,$
by $\delta\leq\varepsilon$, we get
(3.10) $\mathcal{S}\lesssim\varepsilon^{(p-D)/p}+N_{d}\,\varepsilon^{(1-d/d^{\prime})\alpha p}+2\max\Big\{c_{u,0}\,\varepsilon,\;N_{d}\,\varepsilon^{(1-d/d^{\prime})\alpha p+(1/\theta-1)d}\Big\}.$
Note that the comparability constant above, say $C(\lesssim)$, only depends on uniformly fixed constants, and does not depend on $\delta$ or $\epsilon$.
Thus, for any $\varepsilon_{Y}>0$, there is $\varepsilon>0$ small enough such that the right-hand-side of (3.10) times $C(\lesssim)$ is less than $\varepsilon_{Y}$. Fix $\delta_{\varepsilon}$ small enough for all the above assumptions to hold, so that there is $\delta_{\varepsilon_{Y}}^{\prime}=\delta_{\varepsilon}^{d/D^{\prime}}\in(0,1)$ such that for every $\delta_{Y}=\delta^{d/D}\leq\delta_{\varepsilon_{Y}}^{\prime}$ there is a cover $\mathcal{U}$ of $f(E)$, resulting from the corresponding graph of the source $\mathcal{G}(\mathcal{Q_{\delta}})$, with $\delta_{Y}^{1/\theta}\leq|U|\leq\delta$ for all $U\in\mathcal{U}$, and
$\sum_{U\in\mathcal{U}}|U|^{D}<\varepsilon_{Y}.$
This implies $\dim_{\theta}f(E)\leq D$, and the proof is complete by taking $d^{\prime}\rightarrow d_{E}$. ∎
Remark 3.2.
  1. (i)
    We emphasize that the approach we took for the sum over blue cubes with ancestors in $V$, i.e. over $V_{B}^{\prime}$, would not work on the sum over blue cubes $V_{B}^{\text{orig}}$ in $\mathcal{Q}_{\delta}$. That is because the bound on their number from (3.1), (3.2), the bound on the diameters of images by $\delta_{Y}$, and (3.3) would yield
    $\sum_{Q\in V_{B}^{\text{orig}}}|f(Q)|^{D}\leq c_{u,0}\varepsilon\delta^{-d/\theta}\delta_{Y}^{D}=c_{u,0}\varepsilon\delta^{-d/\theta}\delta^{d},$
    with the right-hand side being potentially very large for small $\delta$.
  2. (ii)
    Comparing our terminology to that of Kaufman, red and blue cubes correspond to “major” and “minor” cubes in [54]. Due to the nature of the intermediate dimension, requiring a lower bound on the diameters of covering sets, we had to introduce the class of green cubes to account for sets in the target that are not too large, but are too small. This is not necessary when investigating the distortion of the upper box-counting dimension, and thus a similar notion was not needed in [54].
4. Intermediate dimensions under fractionally smooth Sobolev mappings
4.1. Morrey embedding theorem for fractional Hajłasz–Sobolev, Hajłasz–Triebel–Lizorkin, and Hajłasz–Besov spaces
In order to show that continuous mappings with a finite Hajłasz–Triebel–Lizorkin or Hajłasz–Besov semi-norm are compactly Hölder, we will employ the Morrey embedding theorem.
We begin by establishing such an embedding theorem for Hajłasz–Sobolev spaces $\dot{M}^{s,p}$. To facilitate the formulation of the result, we introduce the following piece of notation: Given $Q,b\in(0,\infty)$, $\sigma\in[1,\infty)$, and a ball $B_{0}\subset X$ of radius $R_{0}\in(0,\infty)$, the measure $\mu$ is said to satisfy the $V(\sigma B_{0},Q,b)$ condition [^1] This condition is a slight variation of the one in [39, p. 197]., provided that, for any $x\in X$ and $r\in(0,\sigma R_{0}]$ satisfying $B(x,r)\subset\sigma B_{0}$,
(4.1) $\mu\left(B(x,r)\right)\geq br^{Q}.$
Note that both locally $Q$-homogeneous and $Q$-Ahlfors regular measures satisfy the $V(\sigma B_{0},Q,b)$ condition for all balls $B_{0}$ having sufficiently small measure.
Theorem 4.1.
Let $(X,d_{X},\mu)$ be a metric measure space and $(Y,d_{Y})$ be a metric space. Let $\sigma\in(1,\infty)$, $B_{0}\subset X$ be a ball of radius $R_{0}\in(0,\infty)$, and suppose $s,p\in(0,\infty)$ satisfy $p>Q/s$. Assume that the measure $\mu$ satisfies the $V(\sigma B_{0},Q,b)$ condition for some $Q,b\in(0,\infty)$, and suppose $u\in\dot{M}^{s,p}(\sigma B_{0}:Y)$. Then, there exists a set $N\subset X$ with $\mu(N)=0$, such that, for all $x,y\in B_{0}\setminus N$,
(4.2) $d_{Y}(u(x),u(y))\leq Cb^{-1/p}[d_{X}(x,y)]^{s-Q/p}\|u\|_{\dot{M}^{s,p}(\sigma B_{0}:Y)}$
where $C$ is a positive constant depending only on $d_{X}$, $s$, $p$, $Q$, and $\sigma$. In particular, if $(Y,d_{Y})$ is complete then $u$ has a Hölder continuous representative of order sQ/ps-Q/p on $B_{0}$, denoted by $u$, satisfying (4.2) for all $x,y\in B_{0}$.
While the proof of Theorem 4.1 is quite similar to the proof of [7, Theorem 3.1] and [3, Theorem 6] for $Y=\mathbb{R}$, there were certain adjustments needed to properly establish the result for mappings with an arbitrary metric space $(Y,d_{Y})$ as a target.
Proof.
Choose $g\in\mathcal{D}^{s}(u)\cap L^{p}(\sigma B_{0})$ such that $\|g\|_{L^{p}(\sigma B_{0})}\approx\|u\|_{\dot{M}^{s,p}(\sigma B_{0}:Y)}$. Without loss of generality, we may assume $\int_{\sigma B_{0}}g^{p}\,d\mu>0$. Indeed, if this integral equals zero, then $g=0$ $\mu$-almost everywhere in $\sigma B_{0}$ which, in turn, implies that $u$ is a constant function $\mu$-almost everywhere in $B_{0}$ and the result is obvious in this scenario.
[…]
Let $N:=E\cup\{x\in\sigma B_{0}:g(x)=\infty\mbox{ or }g(x)=0\}$, where $E\subset\sigma B_{0}$ is a set of measure zero such that the pointwise inequality (2.12) holds for every $x,y\in\sigma B_{0}\setminus E$. Then $N$ is a measurable set satisfying $\mu(N)=0$.
To prove (4.2), we will first show that there exists a point $\xi_{0}\in\sigma B_{0}\setminus N$ such that, for every $x\in B_{0}\setminus N$,
(4.4) $d_{Y}(u(x),u(\xi_{0}))\leq Cb^{-1/p}R_{0}^{s-Q/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p},$
where $C$ is a positive constant depending only on $d$, $s$, $p$, $Q$, and $\sigma$. To this end, for any $k\in\mathbb{Z}$, let
$E_{k}:=\left\{x\in\sigma B_{0}\setminus N:\,g(x)\leq 2^{k}\right\}.$
Clearly $E_{k-1}\subset E_{k}$ for any $k\in\mathbb{Z}$. By (4.3), we find that
(4.5) $\bigcup_{k\in\mathbb{Z}}\left[E_{k}\setminus E_{k-1}\right]=\sigma B_{0}\setminus N.$
It follows from the pointwise inequality (2.12) that $u$ restricted to $E_{k}$ is $2^{k+1}$-Hölder continuous of order $s$, that is,
(4.6) $d_{Y}(u(x),u(y))\leq 2^{k+1}[d_{X}(x,y)]^{s},\ \forall\,x,y\in E_{k}.$
Also, applying the Chebyshev inequality, we have
(4.7) $\mu(\sigma B_{0}\setminus E_{k})=\mu\left(\left\{x\in\sigma B_{0}:\,g(x)>2^{k}\right\}\right)\leq 2^{-kp}\int_{\sigma B_{0}}g^{p}\,d\mu.$
For any $k\in\mathbb{Z}$ and $\gamma\in Y$, let
(4.10) $a_{k}:=\left\{\begin{array}[]{ll}\sup\limits_{E_{k}\cap B_{0}}d_{Y}(u,\gamma)\quad&\text{if }\,E_{k}\cap B_{0}\neq\emptyset,\\[15.0pt] \qquad 0&\text{otherwise.}\end{array}\right$.
Clearly, $a_{k}\leq a_{k+1}$ for any $k\in\mathbb{Z}$.
We will need the following elementary result from [3, Lemma 8].
Lemma 4.2.
For any $x\in X$ and $r\in(0,\infty)$, if $B(x,r)\subset\sigma B_{0}$ and $\mu(B(x,r))\geq 2\mu(\sigma B_{0}\setminus E_{k})$ for some $k\in\mathbb{Z}$, then
$\mu(B(x,r)\cap E_{k})\geq\frac{1}{2}\mu(B(x,r))>0.$
Moving on, let $k_{0}$ be the least integer such that
$2^{k_{0}}\geq\left[\frac{2^{1/Q}}{(\sigma-1)(1-2^{-p/Q})}\right]^{Q/p}\left(bR_{0}^{Q}\right)^{-1/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p}$
or, equivalently,
(4.11) $2^{-k_{0}p/Q}\,\frac{2^{1/Q}b^{-1/Q}}{\left(1-2^{-p/Q}\right)}\,\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/Q}\leq(\sigma-1)R_{0}.$
Clearly,
(4.12) $2^{k_{0}}\approx\left(bR_{0}^{Q}\right)^{-1/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p},$
where the positive equivalence constants depend on $Q$, $p$, $\sigma$, and $d$. The following lemma is a straightfoward modification of [3, Lemma 9].
Lemma 4.3.
Under the above assumptions, one has $\mu(E_{k_{0}})\geq\mu(\sigma B_{0})/2$.
Proof.
Suppose to the contrary that $\mu(E_{k_{0}})<\mu(\sigma B_{0})/2$. Then
(4.13) $\mu(\sigma B_{0}\setminus E_{k_{0}})>\mu(\sigma B_{0})/2.$
By (4.7) and (4.11), we find that
$r:=$  $\,2^{1/Q}b^{-1/Q}[\mu(\sigma B_{0}\setminus E_{k_{0}})]^{1/Q}$
$\leq$  $\,2^{1/Q}b^{-1/Q}2^{-k_{0}p/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/Q}$
$\leq$ $\,(\sigma-1)(1-2^{-p/Q})R_{0}<(\sigma-1)R_{0}.$
Therefore, if $z_{0}\in X$ is the center of the ball $B_{0}$, then $B(z_{0},r)\subset\sigma B_{0}$, so the $V(\sigma B_{0},Q,b)$ condition and (4.13) give
$\mu(\sigma B_{0})\geq\mu(B(z_{0},r))\geq br^{Q}=2\mu(\sigma B_{0}\setminus E_{k_{0}})>\mu(\sigma B_{0}),$
which is an obvious contradiction. This finishes the proof of Lemma 4.3. ∎
Now we prove that, for any integer $k>k_{0}$,
(4.14) $a_{k}\lesssim b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}\,\sum_{j=k_{0}}^{k-1}2^{j(1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(u,\gamma).$
To see this, let $k\in\mathbb{Z}$ satisfy $k>k_{0}$. Observe that, if $E_{k}\cap B_{0}=\emptyset$, then $a_{k}=0$ due to its definition and (4.14) is trivially true. So we only need to consider the case $E_{k}\cap B_{0}\neq\emptyset$ below. In this case, $a_{k}=\sup_{E_{k}\cap B_{0}}d_{Y}(u,\gamma)$. Now, for any $i\in\{0,1,\ldots,k-k_{0}-1\}$, define
$r_{k-i}:=2^{1/Q}b^{-1/Q}2^{-[k-(i+1)]p/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/Q}.$
Then, by (4.11), we conclude that
$r_{k}+r_{k-1}+\ldots+r_{k_{0}+1}$
$\quad=2^{1/Q}b^{-1/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/Q}\left(\sum_{i=0}^{k-k_{0}-1}2^{-[k-(i+1)]p/Q}\right)$
(4.15) $\quad<2^{-k_{0}p/Q}\frac{2^{1/Q}b^{-1/Q}}{\left(1-2^{-p/Q}\right)}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/Q}\leq(\sigma-1)R_{0}.$
The importance of (4.1) will reveal itself shortly. Since $E_{k}\cap B_{0}\neq\emptyset$, we can choose a point $x_{k}\in E_{k}\cap B_{0}$ arbitrarily. We now use induction with respect to $i$ to define a sequence $x_{k-i}\in\sigma B_{0}$, $i\in\{1,\ldots,k-k_{0}\}$, such that $x_{k-1}\in E_{k-1}\cap B(x_{k},r_{k})$,
$x_{k-2}\in E_{k-2}\cap B(x_{k-1},r_{k-1}),\ \ldots,\ x_{k_{0}}\in E_{k_{0}}\cap B(x_{k_{0}+1},r_{k_{0}+1}).$
To be precise, for $i=1$, we choose $x_{k-1}$ as follows. First, by (4.1), we find that, for any point $y\in B(x_{k},r_{k})$,
$d_{X}(z_{0},y)\leq d_{X}(z_{0},x_{k})+d_{X}(x_{k},y)<R_{0}+r_{k}<R_{0}+(\sigma-1)R_{0}=\sigma R_{0},$
which implies $B(x_{k},r_{k})\subset\sigma B_{0}$. As such, applying the $V(\sigma B_{0},Q,b)$ condition and (4.7), we conclude that
$\mu(B(x_{k},r_{k}))\geq br_{k}^{Q}=2\;2^{-(k-1)p}\int_{\sigma B_{0}}g^{p}\,d\mu\geq 2\mu(\sigma B_{0}\setminus E_{k-1}),$
which, together with Lemma 4.2, further implies that $\mu(E_{k-1}\cap B(x_{k},r_{k}))>0$ and hence we can find a point $x_{k-1}\in E_{k-1}\cap B(x_{k},r_{k})$. Clearly $x_{k-1}\in\sigma B_{0}$. If $k-k_{0}=1$, then we are done, so suppose that $k-k_{0}>1$ and assume that we already selected points $x_{k-1},\ldots,x_{k-i}$ for some $1\leq i<k-k_{0}$ satisfying
$x_{k-j}\in\sigma B_{0}\cap E_{k-j}\cap B(x_{k-j+1},r_{k-j+1})\ \mbox{ for\ any\ $j\in\{1,\ldots,i\}$}.$
It remains to select
$x_{k-(i+1)}\in\sigma B_{0}\cap E_{k-(i+1)}\cap B(x_{k-i},r_{k-i}).$
By (4.1), we find that, for any $y\in B(x_{k-i},r_{k-i})$,
$d_{X}(y,x_{k})$ $\leq d_{X}(y,x_{k-i})+d(x_{k-i},x_{k-i+1})+\ldots+d_{X}(x_{k-1},x_{k})$
$<r_{k-i}+r_{k-i+1}+\ldots+r_{k}\leq(\sigma-1)R_{0},$
which, together with $x_{k}\in B_{0}$, further implies that
$d_{X}(z_{0},y)$ $\leq d_{X}(z_{0},x_{k})+d_{X}(x_{k},y)<R_{0}+(\sigma-1)R_{0}=\sigma R_{0}.$
Thus, $B(x_{k-i},r_{k-i})\subset\sigma B_{0}$, and then the $V(\sigma B_{0},Q,b)$ condition and (4.7) imply
$\mu(B(x_{k-i},r_{k-i}))$ $\geq br_{k-i}^{Q}$
$=2\;2^{-[k-(i+1)]p}\int_{\sigma B_{0}}g^{p}\,d\mu\geq 2\mu(\sigma B_{0}\setminus E_{k-(i+1)}).$
Applying Lemma 4.2, we have $\mu(E_{k-(i+1)}\cap B(x_{k-i},r_{k-i}))>0$ and can find a point
$x_{k-(i+1)}\in E_{k-(i+1)}\cap B(x_{k-i},r_{k-i}).$
Clearly $x_{k-(i+1)}\in\sigma B_{0}$. That finishes the inductive argument on the choose of $\{x_{k_{0}},\ldots,x_{k-1}\}$.
Notice that, for any $i\in\{0,1,\ldots,k-k_{0}-1\}$,
$d_{X}(x_{k-i},x_{k-(i+1)})<r_{k-i}$ $=2^{1/Q}b^{-1/Q}2^{-[k-(i+1)]p/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/Q}.$
Also, recall that, by (4.6), $u$ restricted to $E_{k-i}$ is $2^{k-i+1}$-Hölder continuous of order $s$. From these and the fact that $x_{k-i},x_{k-(i+1)}\in E_{k-i}$ and $x_{k_{0}}\in E_{k_{0}}$, it follows that
$d_{Y}(u(x_{k}),\gamma)$ $\leq\sum_{i=0}^{k-k_{0}-1}d_{Y}(u(x_{k-i}),u(x_{k-(i+1)}))+d_{Y}(u(x_{k_{0}}),\gamma)$
$\leq\sum_{i=0}^{k-k_{0}-1}2^{k-i+1}[d_{X}(x_{k-i},x_{k-(i+1)})]^{s}+\sup_{E_{k_{0}}}d_{Y}(u,\gamma)$
$\lesssim 2^{s/Q}b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}\,\sum_{i=0}^{k-k_{0}-1}2^{[k-(i+1)](1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(u,\gamma)$
(4.16) $\lesssim b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}\,\sum_{j=k_{0}}^{k-1}2^{j(1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(u,\gamma).$
Since $x_{k}\in E_{k}\cap B_{0}$ was selected arbitrarily, taking the supremum in (4.1) over all $x_{k}\in E_{k}\cap B_{0}$, we obtain the desired estimate in (4.14).
Observe that, on the one hand, for any $x,y\in\sigma B_{0}$,
(4.17) $|\sigma B_{0}|\leq 2\sigma R_{0}.$
On the other hand, by Lemma 4.3 we have $\mu(E_{k_{0}})>0$, and so we can choose a point $\xi_{0}\in E_{k_{0}}\subset\sigma B_{0}\setminus N$. Taking $\gamma=u(\xi_{0})$ and applying the Hölder continuity (4.6), gives
$\sup_{E_{k_{0}}}d_{Y}(u,u(\xi_{0}))$ $\leq 2^{k_{0}+1}|\sigma B_{0}|^{s}\leq 2^{k_{0}+1}\left[2\sigma R_{0}\right]^{s}$
(4.18) $\lesssim R_{0}^{s}\left(bR_{0}^{Q}\right)^{-1/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p},$
where the implicit positive constant depends only on $Q$, $p$, $\sigma$, $s$, and $d$.
With $\gamma=u(\xi_{0})$ as in (4.1) and $\{a_{k}\}_{k\in\mathbb{Z}}$ as in (4.10), observe that since $2^{1-sp/Q}<1$, by (4.14), we find that, for any integer $k>k_{0}$,
(4.19) $a_{k}\lesssim b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}2^{k_{0}(1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(u,u(\xi_{0})).$
However, since, for every $k\leq k_{0}$,
$a_{k}:=\sup_{E_{k}\cap B_{0}}d_{Y}(u,u(\xi_{0}))\leq\sup_{E_{k_{0}}}d_{Y}(u,u(\xi_{0})),$
we actually have that (4.19) holds for all $k\in\mathbb{Z}$. From this, (4.12), and (4.1), we deduce that, for every $k\in\mathbb{Z}$,
(4.20) $a_{k}:=\sup_{E_{k}\cap B_{0}}d_{Y}(u,u(\xi_{0}))\lesssim b^{-1/p}R_{0}^{s-Q/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p}.$
Noticing that the right-hand side of (4.20) is a positive constant independent of $k$, by (4.5) and the definition of $a_{k}$ in (4.10), we conclude that the function $d_{Y}\left(u,u(\xi_{0})\right)$ is bounded on $B_{0}\setminus N$ by the right-hand side of (4.20) (modulo a positive constant). This proves (4.4).
Next we show the Hölder continuity of $u$ along with the estimate (4.2). To this end, fix $x,y\in B_{0}\setminus N$. If $2d_{X}(x,y)\leq(\sigma-1)R_{0}/\sigma$, let $R_{1}:=2d_{X}(x,y)$. Clearly, $x,y\in B_{1}:=B(x,R_{1})$. Moreover, if $z_{0}$ denotes the center of $B_{0}$, then, for any $z\in\sigma B_{1}$, we have
$d_{X}(z_{0},z)$ $\leq d_{X}(z_{0},x)+d_{X}(x,z)<R_{0}+\sigma R_{1}<R_{0}+(\sigma-1)R_{0}=\sigma R_{0},$
which implies that $\sigma B_{1}\subset\sigma B_{0}$. From this and the assumption that $\mu$ satisfies the $V(\sigma B_{0},Q,b)$ condition, it follows that $\mu$ also satisfies the $V(\sigma B_{1},Q,b)$ condition and therefore, by the estimate (4.4) applied to $B_{1}$ in place of $B_{0}$, we can find a point $\xi_{1}\in\sigma B_{1}\setminus N$ such that,
$d_{Y}(u(x),u(y))$ $\leq d_{Y}(u(x),u(\xi_{1}))+d_{Y}(u(\xi_{1}),u(y))$
$\lesssim b^{-1/p}R_{1}^{s-Q/p}\,\left(\,\int_{\sigma B_{1}}g^{p}\,d\mu\right)^{1/p}$
$\approx b^{-1/p}[d_{X}(x,y)]^{s-Q/p}\left(\,\int_{\sigma B_{1}}g^{p}\,d\mu\right)^{1/p}$
$\lesssim b^{-1/p}[d_{X}(x,y)]^{s-Q/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p}.$
If $2d_{X}(x,y)>(\sigma-1)R_{0}/\sigma$, then (4.4) immediately gives
$d_{Y}(u(x),u(y))$ $\leq d_{Y}(u(x),u(\xi_{0}))+d_{Y}(u(\xi_{0}),u(y))$
$\lesssim b^{-1/p}R_{0}^{s-Q/p}\,\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p}$
$\lesssim b^{-1/p}[d_{X}(x,y)]^{s-Q/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p}.$
These estimates imply that (4.2) holds for every $x,y\in B_{0}\setminus N$. Since $B_{0}\setminus N$ is dense in $B_{0}$ [here we are using the fact that $\mu$ is positive and finite on balls], we can rely on the completeness of $(Y,d_{Y})$ to extend $u$ to a Hölder continuous function of order sQ/ps-Q/p on $B_{0}$, that satisfies (4.2) for every $x,y\in B_{0}$. This completes the proof of Theorem 4.1. ∎
Remark 4.4.
From the proof of Theorem 4.1, if $u\in\dot{M}^{s,p}(X:Y)$ then the null set $N$, and hence also the Hölder continuous representative of $u$, can be chosen independent of the ball $B_{0}$.
We now establish the Morrey embedding theorem for Hajłasz–Triebel–Lizorkin and Hajłasz–Besov on locally $Q$-homogeneous metric measure spaces. We will need the following result, which was proven for real-valued functions in [7, Proposition 2.4]. The same proof is also valid for metric space-valued functions and therefore we omit the details.
Lemma 4.5.
Let $(X,d_{Y},\mu)$ be a metric space equipped with a nonnegative Borel measure $\mu$, $(Y,d_{Y})$ be a metric space, and suppose that $s,p\in(0,\infty)$ and $q\in(0,\infty]$. Then
  1. (1)
    $\dot{M}^{s}_{p,q}(X:Y)\hookrightarrow\dot{M}^{s}_{p,\infty}(X:Y)$;
  2. (2)
    $\dot{M}^{s}_{p,\infty}(X:Y)=\dot{M}^{s,p}(X:Y)$ as sets, with equal semi-norms;
  3. (3)
    if $q\in(0,p]$, then $\dot{N}^{s}_{p,q}(X:Y)\hookrightarrow\dot{M}^{s,p}(X:Y)\hookrightarrow\dot{N}^{s}_{p,\infty}(X:Y)$;
  4. (4)
    for any $\varepsilon\in(0,s)$, there exists a constant $C\in(0,\infty)$ such that, if $B\subset X$ is a ball with radius $r\in(0,\infty)$, then $\dot{N}^{s}_{p,q}(B:Y)\hookrightarrow\dot{M}^{\varepsilon,p}(B:Y)$, where $\|u\|_{\dot{M}^{\varepsilon,p}(B:Y)}\leq Cr^{s-\varepsilon}\|u\|_{\dot{N}^{s}_{p,q}(B:Y)}$ for all $u\in\dot{N}^{s}_{p,q}(B:Y)$.
Corollary 4.6.
Let $(X,d,\mu)$ be a metric measure space, where $\mu$ is locally $Q$-homogeneous for some $Q\in(0,\infty)$, and let $(Y,d_{Y})$ be a metric space. Let $s\in(0,\infty)$, $p\in(Q/s,\infty)$, and $q\in(0,\infty]$, and assume $f:X\to Y$ is a continuous function such that $\|f\|_{\dot{M}^{s}_{p,q}(B:Y)}$ is finite for each fixed ball $B\subset X$. Then, for any compact set $K\subset X$, there exist constants $C_{K}\in[1,\infty)$ and $R_{K}\in(0,\infty)$, both of which are independent of $f$, such that, for all balls $B_{0}:=B(x_{0},R_{0})$ with $x_{0}\in K$ and $R_{0}\in(0,R_{K})$, and one has
(4.21) $|f(x)-f(y)|\leq C_{K}[d(x,y)]^{s-Q/p}\frac{R_{0}^{Q/p}}{[\mu(2B_{0})]^{1/p}}\,\|f\|_{\dot{M}^{s}_{p,q}(2B_{0}:Y)},$
for all $x,y\in B_{0}$. The above statement is also valid with the Hajłasz–Triebel–Lizorkin space $\dot{M}^{s}_{p,q}$ replaced by the Hajłasz–Besov space $\dot{N}^{s}_{p,q}$.
Proof.
Fix a compact set $K\subset X$. We claim that there exists a constant $\kappa\in(0,\infty)$ such that
(4.22) $\kappa\left(\frac{r_{1}}{r_{2}}\right)^{Q}\leq\frac{\mu(B(y,r_{1}))}{\mu(B(x,r_{2}))},$
for all $x,y\in K$ and $0<r_{1}<r_{2}<\tilde{R}_{\text{hom}}(K)/2$ satisfying $B(y,r_{1})\subset B(x,r_{2})$, where $\tilde{R}_{\text{hom}}(K)\in(0,\infty)$ is as in the $Q$-homogeneous condition (2.14). Suppose $x,y\in K$ and $0<r_{1}<r_{2}<\tilde{R}_{\text{hom}}(K)/2$ are such that $B(y,r_{1})\subset B(x,r_{2})$. Since $0<r_{1}<2r_{2}<\tilde{R}_{\text{hom}}(K)$, by (2.14) we have
$\frac{\mu(B(y,2r_{2}))}{\mu(B(y,r_{1}))}\leq\tilde{C}_{\text{hom}}(K)\left(\frac{2r_{2}}{r_{1}}\right)^{Q},$
which, combined with the observation $B(x,r_{2})\subset B(y,2r_{2})$, gives
$\mu(B(x,r_{2}))\leq\mu(B(y,2r_{2}))\leq\tilde{C}_{\text{hom}}(K)\left(\frac{2r_{2}}{r_{1}}\right)^{Q}\mu(B(y,r_{1})).$
Thus, (4.22) holds with $\kappa:=[2^{Q}\tilde{C}_{\text{hom}}(K)]^{-1}$.
Moving on, let $R_{K}:=\tilde{R}_{\text{hom}}(K)/4$ and $B_{0}:=B(x_{0},R_{0})$ a ball with $x_{0}\in K$ and $R_{0}\in(0,R_{K})$. Then (4.22) applied with $B(x,r_{2}):=B(x_{0},2R_{0})=2B_{0}$ implies that the measure $\mu$ satisfies the $V(2B_{0},Q,b)$ condition with $b:=\kappa\mu(2B_{0})(2R_{0})^{-Q}$. On the other hand, since $\|f\|_{\dot{M}^{s}_{p,q}(2B_{0}:Y)}$ is finite, it follows from (1) and (2) in Lemma 4.5 that $\|f\|_{\dot{M}^{s,p}(2B_{0}:Y)}$ is also finite. Combining these observations with Theorem 4.1, we conclude that there exists a set $N\subset X$ with $\mu(N)=0$, such that the inequality in (4.2) holds for all $x,y\in B_{0}\setminus N$. However, since $f$ is continuous and $B_{0}\setminus N$ is dense in $B_{0}$ [here we are using the fact that $\mu$ is positive and finite on balls], we actually have that (4.2) holds for all $x,y\in B_{0}$. This completes the proof of (4.21) for $\dot{M}^{s}_{p,q}$ spaces,
To obtain (4.21) for $\dot{N}^{s}_{p,q}$ spaces, choose $\varepsilon\in(0,s)$ close enough to $s$ so that $p\in(Q/\varepsilon,\infty)$. By Lemma 4.5(4), $\dot{N}^{s}_{p,q}(2B_{0}:Y)\hookrightarrow\dot{M}^{\varepsilon,p}(2B_{0}:Y)$ with
(4.23) $\|f\|_{\dot{M}^{\varepsilon,p}(2B_{0}:Y)}\leq CR_{0}^{s-\varepsilon}\|f\|_{\dot{N}^{s}_{p,q}(2B_{0}:Y)},$
for some constant $C\in(0,\infty)$ that is independent of $f$ and the ball $B_{0}$. Given this and the fact that $p\in(Q/\varepsilon,\infty)$, by arguing as we did in the proof of (4.4) with the function $f\in\dot{M}^{\varepsilon,p}(\sigma B_{0}:Y)$, we conclude that there exist a set $N\subset X$ (which can be chosen independent of $B_{0}$) with $\mu(N)=0$ and a point $\xi_{0}\in 2B_{0}\setminus N$ such that, for every $x\in B_{0}\setminus N$,
$d_{Y}(f(x),f(\xi_{0}))$ $\leq b^{-1/p}R_{0}^{\varepsilon-Q/p}\|f\|_{\dot{M}^{\varepsilon,p}(2B_{0}:Y)}$
(4.24) $\lesssim b^{-1/p}R_{0}^{s-Q/p}\|f\|_{\dot{N}^{s}_{p,q}(2B_{0}:Y)}.$
where the implicit constant depends only on $d$, $s$, $\varepsilon$, $p$, and $Q$. With (4.1) in hand, an argument analogous to the one used at the end Theorem 4.1 (keeping in mind that $b:=\kappa\mu(2B_{0})(2R_{0})^{-Q}$) gives that the inequality in (4.21) holds pointwise almost everywhere in $B_{0}$. Now, proceeding as in the case of $\dot{M}^{s}_{p,q}$ spaces, we conclude that (4.21) in fact holds pointwise everywhere in $B_{0}$. This finishes the proof of Corollary 4.6. ∎
4.2. The proof of Theorem 1.3
To prove Theorem 1.3, we require one final technical lemma, whose statement relies on the following notion: Given a metric measure space $(X,d_{X},\mu)$ and a threshold $R\in(0,\infty)$, recall that the restricted Hardy–Littlewood maximal function of $f\in L^{1}_{{\scriptstyle{loc}}}(X)$ is defined by setting, for each $x\in X$,
[…]
Suppose $\mu$ is locally $Q$-homogeneous for some $Q\in(0,\infty)$ and let $\tilde{R}_{\text{hom}}\in(0,\infty)$ be as in the local $Q$-homogeneity condition (2.14). Then for every non-empty compact set $K\subseteq X$ and $p\in(1,\infty)$, there exists a constant $C^{\prime}\in(0,\infty)$ such that,
(4.25) $\|M_{R}f\|_{L^{p}(K)}\leq C^{\prime}\|f\|_{L^{p}(K^{\prime})},$
for all $R\in(0,\tilde{R}_{\text{hom}}(K)/10)$ and $f\in L^{p}(X)$, where $K^{\prime}\subseteq X$ is any measurable set that contains every ball centered in $K$ with radius at most $\tilde{R}_{\text{hom}}/10$. Indeed, this follows from arguing as in the proof of the Maximal Function Theorem, see, for instance, [44, Chapter 2] or [47, Theorem 3.5.6].
We can now state the lemma alluded to above.
Lemma A.
Let $(X,d_{X},\mu)$ be a metric measure space, where $\mu$ is locally doubling, and fix $t\in[1,p)$ and $\tau\in(0,1)$. For each compact set $K\subset X$ there is a constant $C_{K}^{\prime}\geq 1$ and a radius $R_{K}^{\prime}>0$ such that for all nonnegative functions $g\in L^{p}(X)$,
[…]
for all $x\in K$ and $0<r<R_{K}^{\prime}$.
Lemma A is a corollary of the Maximal Function Theorem, see [44, Chapter 2] and [12, Lemma 3.3] for a proof.
Suppose $f:X\rightarrow Y$ is continuous with a finite Hajłasz–Triebel–Lizorkin semi-norm $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}$, or finite a Hajłasz–Besov semi-norm $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}$. Note that for every $t\in[1,p)$ and every ball $B\subset X$ we have $\|f\|_{\dot{M}^{s}_{t,q}(B:Y)}$ or $\|f\|_{\dot{N}^{s}_{t,q}(B:Y)}$ is finite. We plan on using Corollary 4.6 and Lemma A to show that $f$ belongs to the appropriate compactly Hölder class, which is enough to achieve (1.4) and (1.5).
Proof of Theorem 1.3.
Let $E\subset X$ be compact and $\varepsilon\in(0,1)$. Note that the centers of balls that cover $E$ in the definition of compactly Hölder mappings do not necessarily lie in $E$, while it is a requirement for the inequalities in Corollary 4.6 and Lemma A. Thus, we need to apply these properties to a potentially larger compact set than $E$. Take any point $x_{E}\in E$. Then a straightforward calculation shows that
(4.27) $E\cup\left(\bigcup_{x\in E}B(x,1/4)\right)\subset B(x_{E},|E|+1/2).$
Thus, if $K$ is the closure of $B(x_{E},|E|+1/2)$, and $\mathcal{B}$ is a covering of $E$ by balls of radius at most $1/10$, then all elements of $\mathcal{B}$ lie entirely in $K$, along with their centers. Our plan is to apply Corollary 4.6 and Lemma A with the set $K$, which is compact by virtue of $X$ being proper. With this goal in mind, consider the radius
(4.28) $r_{E}:=\min\left\{\frac{R_{K}}{2},\frac{R_{K}^{\prime}}{2},\frac{R_{\text{hom}}(K)}{10},\frac{1}{10}\right\},$
where $R_{K},R_{K}^{\prime}$ are as in Corollary 4.6 and Lemma A, respectively, and $R_{\text{hom}}(K):=\tilde{R}_{\text{hom}}(K)/3$ with $\tilde{R}_{\text{hom}}(K)$ as in the local $Q$-homogeneity condition (2.14) for the compact set $K$.
Suppose $\{B(x_{i},r)\}_{i\in I}$ is a cover of $E$ with $r<r_{E}$ and $B(x_{i},\varepsilon r)\cap B(x_{j},\varepsilon r)=\emptyset$ for all distinct $i,j\in I$. We will show that an inequality of the form (2.9) holds. Let us first consider the case when $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}$ is finite. By (1) and (2) in Lemma 4.5, we can assume $\|f\|_{\dot{M}^{s,p}(X:Y)}$ is finite. Let $g\in\mathcal{D}^{s}(f)\cap L^{p}(X)$ and fix $t\in(\min\{1,Q/s\},p)$. Then it follows from Hölder’s inequality that $\|f\|_{\dot{M}^{t,p}(B:Y)}$ is finite for each fixed ball $B\subset X$, where the pointwise restriction of $g$ to $B$ serves as an $s$-gradient of $f$. By this, (4.28), and Corollary 4.6, we have that there exists a constant $C_{k}\in[1,\infty)$, which is independent of $f$, such that, for all balls $B_{i}:=B(x_{i},r)$, the following inequality holds
[…]
for all $x,y\in B_{i}$. Setting $\alpha:=s-Q/t$, we deduce from the preceding inequality and the definition of $|f|_{\alpha,B_{i}}$, that
[…]
where $\tilde{g}:=M_{r_{E}}(g^{t})\in L^{p/t}(K)\subset L^{1}(K)$ by (4.25). Since the measure $\mu$ is locally $Q$-homogeneous, by (2.15) and (4.28), the quotient $\frac{|B_{i}|^{Q}}{\mu(\varepsilon B_{i})}$ is bounded from above by $\tilde{C}:=C_{\text{hom}}(K)\left(\frac{1}{\varepsilon}\right)^{Q}$. Hence, due to $B(x_{i},\varepsilon r)\cap B(x_{j},\varepsilon r)=\emptyset$ and
$\sum_{i\in I}\int_{\varepsilon B_{i}}\tilde{g}\,d\mu=\int_{\bigcup\limits_{i\in I}\varepsilon B_{i}}\tilde{g}\,d\mu\leq\int_{K}\tilde{g}\,d\mu<\infty,$
there is $C_{E}=C_{K}^{t}C_{K}^{\prime}\tilde{C}\left(\int_{K}\tilde{g}\,d\mu\right)<\infty$ such that
$\sum\limits_{i\in I}|f|_{\alpha,B_{i}}^{t}\leq\sum\limits_{i\in I}C_{K}^{q}C_{K}^{\prime}\tilde{C}\int_{\varepsilon B_{i}}\tilde{g}\,d\mu\leq C_{E}.$
Since $E$ and $\varepsilon$ were arbitrary, and given that $K$ depends only on $E$, we deduce that $f\in CH^{t,s-Q/t}$ for all $t\in(Q/s,p)$ as wanted. As such, it follows from Theorem 1.1 that
$\dim_{\theta}f(E)\leq\frac{td_{E}(\theta)}{st-Q+d_{E}(\theta)},$
for all $t\in(Q/s,p)$, which, by letting $t\rightarrow p$, gives (1.4).
Suppose that $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}$ is finite and let $\vec{g}:=\{g_{k}\}_{k\in\mathbb{Z}}\in\mathbb{D}^{s}(f)\cap\ell^{q}(L^{p}(X))$. By Lemma 4.5(3), if $q\in(0,p]$, then $\dot{N}^{s}_{p,q}(X:Y)\hookrightarrow\dot{M}^{s,p}(X:Y)$ and the desired conclusions follow from what we have already proven. Thus, suppose $q>p$ and fix $t\in(\min\{1,Q/s\},p)$. Then it follows from Hölder’s inequality that $\|f\|_{\dot{N}^{s}_{t,q}(B:Y)}$ is finite for each fixed ball $B\subset X$, where the pointwise restriction of the functions in the sequence $\{g_{k}\}_{k\in\mathbb{Z}}$ to $B$ serves as a fractional $s$-gradient of $f$. By this, (4.28), and Corollary 4.6, we have that there exists a constant $C_{k}\in[1,\infty)$, which is independent of $f$, such that, for all balls $B_{i}:=B(x_{i},r)$, the following inequality holds
[…]
for all $x,y\in B_{i}$. Arguing as in the case when $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}$ is finite, we have
$|f|_{\alpha,B_{i}}^{q}\leq C_{K}^{q}(C_{K}^{\prime}\tilde{C})^{q/t}\sum_{k\in\mathbb{Z}}\left(\int_{\varepsilon B_{i}}\tilde{g}_{k}\,d\mu\right)^{q/t},$
where $\alpha:=s-Q/t$, $\tilde{g}_{k}:=M_{r_{E}}(g_{k}^{t})\in L^{p/t}(K)\subset L^{1}(K)$, and $\tilde{C}:=C_{\text{hom}}(K)\left(\frac{1}{\varepsilon}\right)^{Q}$ is the same as before. Since $q>p>t$, we can estimate
$\sum_{i\in I}\sum_{k\in\mathbb{Z}}\left(\int_{\varepsilon B_{i}}\tilde{g}_{k}\,d\mu\right)^{q/t}$ $=\sum_{k\in\mathbb{Z}}\sum_{i\in I}\left(\int_{\varepsilon B_{i}}\tilde{g}_{k}\,d\mu\right)^{q/t}\leq\sum_{k\in\mathbb{Z}}\left(\sum_{i\in I}\int_{\varepsilon B_{i}}\tilde{g}_{k}\,d\mu\right)^{q/t}$
(4.29) $=\sum_{k\in\mathbb{Z}}\left(\int_{\bigcup\limits_{i\in I}\varepsilon B_{i}}\tilde{g}_{k}\,d\mu\right)^{q/t}\leq\sum_{k\in\mathbb{Z}}\left(\int_{K}\tilde{g}_{k}\,d\mu\right)^{q/t}.$
On the other hand, by Hölder’s inequality (used with $p/t>1$) and (4.25), we have
$\int_{K}\tilde{g}_{k}\,d\mu\leq[\mu(K)]^{1-t/p}\left(\int_{K}(\tilde{g}_{k})^{p/t}\,d\mu\right)^{t/p}\leq C^{\prime}[\mu(K)]^{1-t/p}\left(\int_{K^{\prime}}g_{k}^{p}\,d\mu\right)^{t/p},$
for some sufficiently large measurable set $K^{\prime}\subset X$ containing $K$ and some constant $C^{\prime}\in(0,\infty)$ that only depends on the doubling constant $\tilde{C}_{\text{hom}}$ from (2.14). This, combined with (4.2) and the fact that (keeping in mind $\vec{g}:=\{g_{k}\}_{k\in\mathbb{Z}}\in\ell^{q}(L^{p}(X))$)
$\sum_{k\in\mathbb{Z}}\left(\int_{K}g_{k}^{p}\,d\mu\right)^{q/p}=\|\vec{g}\|_{\ell^{q}(L^{p}(K))}^{q}\leq\|\vec{g}\|_{\ell^{q}(L^{p}(X))}^{q}<\infty,$
gives
$\sum_{i\in I}\sum_{k\in\mathbb{Z}}\left(\int_{\varepsilon B_{i}}\tilde{g}_{k}\,d\mu\right)^{q/t}\leq(C^{\prime})^{q/t}[\mu(K)]^{q(1-t/p)/t}\sum_{k\in\mathbb{Z}}\left(\int_{K}g_{k}^{p}\,d\mu\right)^{q/p}<\infty.$
Thus, for $C_{E}:=C_{K}^{q}(C_{K}^{\prime}\tilde{C}C^{\prime})^{q/t}[\mu(K)]^{q(1-t/p)/t}\sum_{k\in\mathbb{Z}}\left(\int_{K}g_{k}^{p}\,d\mu\right)^{q/p}<\infty$, we have
$\sum\limits_{i\in I}|f|_{\alpha,B_{i}}^{q}\leq C_{K}^{q}(C_{K}^{\prime}\tilde{C})^{q/t}\sum_{i\in I}\sum_{k\in\mathbb{Z}}\left(\int_{\varepsilon B_{i}}\tilde{g}_{k}\,d\mu\right)^{q/t}\leq C_{E},$
which implies $f\in CH^{q,s-Q/t}$ for all $t\in(Q/s,p)$. As such, it follows from Theorem 1.1 that
$\dim_{\theta}f(E)\leq\frac{qd_{E}(\theta)}{(s-Q/t)q+d_{E}(\theta)}$
for all $t\in(Q/s,p)$, which, by letting for $t\rightarrow p$, implies
$\dim_{\theta}f(E)\leq\frac{qd_{E}(\theta)}{(s-Q/p)q+d_{E}(\theta)}.$
This completes the proof of Theorem 1.3. ∎
5. Hausdorff and Minkowski dimension distortion
We first note that $X$ supporting a $Q$-Poincaré inequality implies that $X$ is connected [47, Proposition 8.1.6] and, thus, uniformly perfect. As a result, Corollary 1.2 (i), (ii) are direct implications of [21, Theorem 1.2] and [46, Theorem 9.3], respectively. See also the proof of [21, Corollary 1.3] for more details.
Suppose $Q>1$ and $(X,d,\mu)$ is a proper, $Q$-homogeneous metric measure space, and $(Y,d_{Y})$ is arbitrary.
Proof of Theorem 1.4.
(i): Let $E\subset X$ non-empty with $d_{E}=\dim_{H}E<Q$. Due to Theorem 1.3, it is enough to prove how a $(p,\alpha)$-compactly Hölder mapping $f:X\rightarrow Y$ distorts the Hausdorff dimension of $E$.
Let $d>d_{E}$ and $D:=\tfrac{pd}{\alpha p+d}$. By the formulation of the Hausdorff dimension using the dyadic cube systems of Theorem A (see [20, Theorem 1.1 (i)]), we have that for every $\varepsilon>0$ and every $\delta\in(0,\varepsilon)$ there is a cover of $E$ by dyadic cubes $\{Q_{i}^{k_{i}}\}_{i\in I}$, with $b^{k_{i}}\leq\delta$ for all $i\in I$, such that
(5.1) $\sum_{i\in I}|Q_{i}^{k_{i}}|<\varepsilon.$
By Theorem A (iii) and (2.11), we have
$|f(Q_{i}^{k_{i}})|\leq|f(B(Q_{i}^{k_{i}}))|\leq C_{i}|B(Q_{i}^{k_{i}})|^{\alpha}\lesssim C_{i}b^{k_{i}\alpha}.$
This implies that
(5.2) $|f(Q_{i}^{k_{i}})|\lesssim C_{E}\delta^{\alpha},$
and that
$\sum_{i\in I}|f(Q_{i}^{k_{i}})|^{D}\lesssim\sum_{i\in I}C_{i}^{D}b^{Dk_{i}\alpha}.$
By an application of Hölder’s inequality for $p/D>1$, and noting that $D\alpha p(p-D)^{-1}=d$, similarly to the proof of Theorem 1.1, we have
$\sum_{i\in I}|f(Q_{i}^{k_{i}})|^{D}\lesssim C_{E}^{D/p}\left(\sum_{i\in I}b^{k_{i}d}\right)^{\tfrac{D-p}{p}}\lesssim\left(\sum_{i\in I}|Q_{i}^{k_{i}}|^{d}\right)^{\tfrac{D-p}{p}}.$
By (5.1) and (5.2), the above relation implies that $\{f(Q_{i}^{k_{i}})\}_{i\in I}$ is an appropriate cover of $f(E)$ to yield $\dim_{H}f(E)\leq D$. Letting $d\rightarrow d_{E}$ finishes the proof.
(ii): This follows directly by [21, Theorem 1.1] and Theorem 1.3. ∎
Remark 5.1.
It should be noted that in the case where $X,Y$ are uniformly perfect, letting $\theta\rightarrow 1$ in (1.1) yields Theorem 1.4 (ii), without the use of [21, Theorem 1.1]. For Theorem 1.4 (i), although the proof is a very simplified case of that of Theorem 1.1 for $\theta=0$, simply letting $\theta\rightarrow 0$ in (1.1) does not generally yield the desired inequality. This is due to the intermediate dimensions not necessarily being continuous at $\theta=0$ (see for instance [28]).